### v1

# WaveCert: Formal Compiler Validation for Asynchronous Dataflow Programs

Abstract.
Coarse-grained reconfigurable arrays (CGRAs), or dataflow architectures in general, have gained attention in recent years due to their promising power efficiency compared to traditional von Neumann architectures. To program these architectures using ordinary languages such as C, a dataflow compiler must transform the original sequential, imperative program into an equivalent dataflow graph, composed of dataflow operators running in parallel. This transformation is challenging since the asynchronous nature of dataflow graphs allows out-of-order execution of operators, leading to behaviors not present in the original imperative programs.
We address this challenge by developing a translation validation technique for dataflow compilers to ensure that the dataflow program has the same behavior as the original imperative program on all possible inputs and schedules of execution. We apply this method to a state-of-the-art dataflow compiler targeting the RipTide CGRA architecture. Our tool uncovers 7 compiler bugs where the compiler outputs incorrect dataflow graphs, including a data race that is otherwise hard to discover via testing. After repairing these bugs, our tool verifies the correct compilation of all programs in the RipTide benchmark suite.
1. Introduction
Even as Moore’s Law ends, many applications still demand better power efficiency, e.g., for high-performance machine learning or for ultra-low-power environmental monitoring. A flood of new hardware architectures have emerged to meet this demand, but one attractive approach that retains general-purpose programmability is a class of dataflow architectures called coarse-grained reconfigurable arrays (CGRAs) (cgra-survey, ). In a CGRA architecture, an array of processing elements (PEs) is physically laid out on the chip. Each PE can be configured to function as a high-level operator that performs an arithmetic operation, a memory operation, or a control-flow operation. These PEs are interconnected via an on-chip network, reducing communication and data movement costs compared to traditional von Neumann architectures, thus making a CGRA more power efficient.
Abstractly, CGRAs run dataflow programs, or networks of operators that execute independently and asynchronously. While CGRAs are able to run dataflow programs efficiently due to their inherent parallelism, this parallelism also introduces new, subtle correctness issues due to the potential for data races. As a result, dataflow programs are not written directly, but are instead compiled from ordinary, imperative languages. However, the gap between imperative and dataflow programming means that these compilers are complex; to maximize performance, they must use a variety of program analyses to convert imperative programs into equivalent distributed ones.
As CGRA architectures continue to emerge and rapidly evolve, we argue that formal verification — mathematically sound proofs of system correctness — should be integrated early on; past experience teaches us that belatedly retrofitting them may be costly or simply infeasible. Since CGRAs cannot execute as intended without correct dataflow programs, we begin this effort by verifying the translation from imperative programs to dataflow programs.
In this work, we propose using translation validation (pnueli-tv, ) to certify the results of dataflow compilers. In our case, translation validation amounts to checking that the output dataflow program from the compiler has the same behaviors as the input imperative program. Compared to traditional testing, translation validation provides a much stronger guarantee, as it checks that the imperative and the dataflow programs behave the same on all possible inputs. Additionally, in contrast to full compiler verification, translation validation allows the use of unverified compilation toolchains, which is crucial for the fast-developing area of CGRA architectures.
Our main challenge for leveraging translation validation is managing the asynchronicity of dataflow programs. Most translation validation works (pnueli-tv, ; necula-tv, ; kundu-tv, ; data-driven-eq, ) prove equivalence between input and output programs via (bi)simulation, or relational invariants between the states of the two programs. Since dataflow programs have exponentially more states than imperative programs (arising from asynchronous scheduling decisions), constructing this bisimulation relation directly would be quite tricky.
Instead, we utilize a two-phase approach. First, we use symbolic execution to prove that there is a bisimulation between the input imperative program and the output dataflow program on a canonical schedule. In the canonical schedule, the dataflow operators are scheduled to fire in a similar order to their counterparts in the original imperative program.
Next, we prove that any possible schedule of the dataflow program must converge to the same final state as the canonical schedule; i.e., it is confluent (term-rewriting, ). For our setting, proving confluence requires showing that the dataflow program does not contain any data races. Taking inspiration from separation logic (separation-logic, ), we augment our symbolic execution with linear permission tokens to track ownership of memory locations. Taken together with the initial bisimulation phase, we can prove that the imperative program is equivalent to the dataflow program on all possible inputs and schedules.
We have implemented this technique in WaveCert  [^1] Available on GitHub at https://github.com/secure-foundations/riptide-verification., a translation validation system targeting the state-of-the-art CGRA architecture RipTide (riptide, ). RipTide features a dataflow compiler from LLVM (llvm, ) programs to dataflow programs. Using our tool, we found 7 compiler bugs where the RipTide dataflow compiler generated incorrect dataflow programs. All of these bugs were confirmed by the developer of the RipTide compiler. One of these bugs allows the compiler to emit a dataflow program with data races, which are hard to discover via testing and costly to fix after deployment.
To summarize, our contributions in this work are:
  • A novel two-phase translation validation technique for certifying compilation results from imperative programs to dataflow programs;
  • An implementation of our technique, WaveCert, which targets the RipTide dataflow compiler; and
  • An evaluation of WaveCert on the RipTide dataflow compiler, which revealed 7 compilation bugs. Each successful verification result takes less than 30 seconds, with most taking around 10 seconds.
Examples of operators.
[…]
2. Preliminaries: CGRAs and RipTide
Dennis et al. (dennis-dataflow, ) introduced early designs for dataflow programming and architectures in the mid 1970s to allow programs to achieve a higher degree of concurrency. In the intervening years, the community has explored a wide variety of dataflow architectures (dataflow-architectures, ; yazdanpanah-dataflow, ). The desire for low-power, general-purpose programmability has recently led to a line of work on coarse-grained reconfigurable arrays (CGRAs) (cgra-survey, ).
To utilize the efficiency and parallelism in CGRAs, the predominant general-purpose way to program CGRAs is to compile from a sequential, imperative program into a dataflow graph via a dataflow compiler. This is used in CGRA projects such as PipeRench (piperench, ), WaveScalar (wavescalar, ), and more recently, RipTide (riptide, ) and Pipestitch (pipestitch, ).
In this work, we focus on the approach taken by RipTide (riptide, ), which has a dataflow compiler from arbitrary LLVM functions to dataflow programs. For an informal description of the semantics of the LLVM intermediate representation (IR), we refer the reader to the official LLVM IR manual (llvm-ir, ). In this work, we use the subset of LLVM IR that is supported by the RipTide compiler, including basic control-flow and integer arithmetic.
The dataflow program output by the RipTide compiler is a Turing-complete representation of programs. It is represented as a dataflow graph in which the nodes are called operators and edges are called channels. Semantically, channels are unbounded FIFO queues used for communication between operators. Operators can be thought of as processes running in parallel with a local state. Repeatedly, each operator: waits to dequeue (a subset of) input channels to get input values; performs a local computation, optionally changing the local state or global memory; and pushes output values to (a subset of) output channels. When an iteration of this loop is done, we say that the operator has executed or fired.
We model the execution of dataflow operators using a fully asynchronous semantics. Different CGRAs would use different scheduling strategies (cgra-survey, ), so we conservatively use asynchronicity to cover all possible schedules.
In Figure 1, we show the five most important types of operators. Reading from left to right, we first have the control flow operators, carry (C) and steer (T). The carry operator is used to support loop variables. It has two local states: in the initial state, it waits for a value from A, outputs it to O, and transitions to the loop state. In the loop state, it waits for values from B and D (decider); if D is true, it outputs O = B; otherwise it discards B and transitions to the initial state.
The steer operator is used for conditional execution. It waits for values from A and D. If D is true, it outputs O = A; otherwise, it discards the value, and outputs nothing.
To utilize the global memory, we have the load (LD) and store (ST) operators. The load operator waits for values from P (base), I (offset), and an optional signal S for memory ordering; reads the memory location P[I]; and outputs the read value to O. Similarly, the store operator waits for values from P (base), I (offset), V (value), and an optional signal S for memory ordering; stores P[I] = V in the global memory; and optionally outputs a finish signal to O.
[…]
To compile an LLVM program to a dataflow program, the RipTide compiler first maps each LLVM instruction to its corresponding dataflow operator, and then it enforces the expected control-flow and memory ordering semantics of the original program. Using a control dependency analysis, it inserts steer operators to selectively enable/disable operators depending on branch/loop conditions. Using a memory ordering analysis, it inserts dataflow dependencies between load/store operators to prevent data races.
These analyses are quite complex and use various optimizations to allow more parallelism than the original LLVM program. As shown in our evaluation in Section 6, this process can easily have bugs. Hence, in our work, we take the translation validation approach to certify the correctness of dataflow compilation.
Figure 2. An example of input/output programs from the RipTide dataflow compiler.
1define void @test(i32* %A, i32* %B, i32 %len) {
2entry:
3  br label %header
4
5header:
6  %i = phi i32 [ 0, %entry ], [ %i_inc, %body ]
7  %cond = icmp slt i32 %i, %len
8  br i1 %cond, label %body, label %end
9
10body:
11  %idx1 = getelementptr i32, i32* %A, i32 %i
12  %A_i = load i32, i32* %idx1
13  %A_i_inc = add i32 %A_i, 1
14  %idx2 = getelementptr i32, i32* %B, i32 %i
15  store i32 %A_i_inc, i32* %idx2
16  %i_inc = add i32 %i, 1
17  br label %header
18
19end:
20  ret void
21}
(a) Input LLVM function.
(b) Output dataflow program.
An example.
Figure 2. An example of input/output programs from the RipTide dataflow compiler.
Figure 3. A trace of execution from the dataflow cut point 2 following the canonical schedule of operators: 1, 2, 4, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2). In addition, steer operators 3 and 5 are also fired in step 3 immediately after they are ready. At the end of the trace, the configuration matches cut point 2 again. For each configuration, we show the state of the channels (solid lines for channels with one value, and dotted lines for empty channels). The lightning symbols indicate which operators are fired in the previous step. Before the firing of an operator, the input channels of the operator should be non-empty (i.e., marked with a solid line). After firing, the input channels are emptied, and the output channels will become non-empty. The exact symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked. At the top of each configuration, we also indicate the list of permission constraints we add at each step.
Cut point 2 trace.
Figure 3. A trace of execution from the dataflow cut point 2 following the canonical schedule of operators: 1, 2, 4, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2). In addition, steer operators 3 and 5 are also fired in step 3 immediately after they are ready. At the end of the trace, the configuration matches cut point 2 again. For each configuration, we show the state of the channels (solid lines for channels with one value, and dotted lines for empty channels). The lightning symbols indicate which operators are fired in the previous step. Before the firing of an operator, the input channels of the operator should be non-empty (i.e., marked with a solid line). After firing, the input channels are emptied, and the output channels will become non-empty. The exact symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked. At the top of each configuration, we also indicate the list of permission constraints we add at each step.
[…]
Index LLVM Dataflow Correspondence
1 %entry Initial config
$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
2
%body
$\to$ %header
See Figure 3
$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
[…]
$\bot$ %end Final config
$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
Cut points for the example.
[…]
3. Overview and an Example
WaveCert performs translation validation on the RipTide dataflow compiler, which compiles LLVM programs to dataflow programs. Given the input program and the compiled output program (along with some hints generated by the compiler), WaveCert performs two checks: first, a bisimulation check, which verifies that the LLVM program is equivalent to the dataflow program on a restricted, canonical schedule of dataflow operators; and second, a confluence check, which proves that the choice of schedule for the dataflow program does not affect the final result.
If a pair of input/output programs passes both checks, the two programs are equivalent. If the bisimulation check fails, there might be a compilation bug that produced an incorrect dataflow. If the bisimulation check succeeds but the confluence check fails, there might be a data race in the output dataflow program.
Example
In Figure 1(b), we have an example input and output program from the dataflow compiler. The input LLVM function @test in Figure 1(a) contains a single loop with loop header block %header and loop body block %body. The loop increments variable %i from 0 to %len, and at each iteration, updates %B[%i] = %A[%i] + 1.
The LLVM function @test gets compiled to the dataflow program in Figure 1(b), in which most operators correspond to an LLVM instruction:
  • The comparison operator 4 corresponds to line 7;
  • The load operator 7 corresponds to line 12;
  • The add operator 8 corresponds to line 13;
  • The store operator 9 corresponds to line 15; and
  • The add operator 6 corresponds to line 16.
The compiler also inserts additional operators to faithfully implement the sequential semantics of the LLVM program. The steer operators marked with T enforce that the operators corresponding to the loop body only execute when the loop condition is true. The carry operator 1 is used for the loop induction variable %i (line 6). The carry operator 2 is used for a loop variable inserted by the compiler to enforce a memory dependency: since arrays A and B could overlap, we have an additional dataflow path from the store operator, passing through the carry operator 2 and the steer operator 5, to the load operator, in order to enforce that the load in the next iteration waits until the previous store has finished.
Bisimulation Check
The first step in WaveCert’s translation validation procedure is to check that the LLVM function is bisimilar to the dataflow program on a canonical schedule of operators, where the operators are executed in the same order as their LLVM instruction counterparts.
To check a bisimulation, we need to construct a relation between the states of the LLVM and dataflow programs. WaveCert constructs this relation by placing corresponding cut points on both the LLVM and the dataflow sides. A cut point symbolically describes a set of pairs of LLVM and dataflow configurations satisfying a particular correspondence condition, and a list of cut points together describes the bisimulation relation. Figure 4 shows the list of cut points for the example. WaveCert places cut point 1 for the initial configurations, cut point 2 for the loop structure, and a final cut point $\bot$ for all final configurations. For each pair of LLVM and dataflow cut points, WaveCert also infers a list of equations expressing the correspondence of symbolic variables in these cut points.
To automatically check that the proposed relation in Figure 4 is indeed a bisimulation, we perform symbolic execution from all pairs of cut points (except for $\bot$) until they reach another pair of cut points. Symbolic execution (symbolic-execution, ) is similar to concrete execution of programs, except with concrete values replaced by symbolic expressions. In this way, we can execute all possible concrete instances of a cut point altogether. On the LLVM side, we first symbolically execute cut point 1, which turns into two symbolic branches: reaching %header after one loop iteration (cut point 2), or failing the loop condition and reaching %end (cut point $\bot$). For LLVM cut point 2, we also have two branches: one reaching cut point 2 again after a loop iteration, the other failing the loop condition and reaching cut point $\bot$.
For the dataflow side, we have a correspondence between a subset of LLVM instructions and a subset of dataflow operators, which is automatically generated by the dataflow compiler. We use this mapping to infer the order in which we fire the dataflow operators, i.e., the canonical schedule. For example, from LLVM cut point 2, there are two branches going to cut point 2 and cut point $\bot$. These two branches have two traces of LLVM instructions (excluding br):
  • For the first branch to cut point 2: lines 6, 7, 11 - 16;
  • For the second branch to cut point $\bot$: lines 6, 7.
For the first branch, the list of LLVM instructions map to dataflow operators 1, 2, 4, 7, 8, 9, 6 (excluding ones that do not have a corresponding operator; operator 2 corresponds to a PHI instruction implicitly inserted during compilation for memory ordering). LLVM instructions in the second branch map to dataflow operators 1, 2, and 4. Therefore, for dataflow cut point 2, we execute these two traces of dataflow operators (and also any steer operator that can be executed), which gives us two dataflow configurations. Figure 3 shows the first branch from cut point 2 to itself. At the end of the execution, we can see that the configuration matches the cut point 2 configuration we started with, except with potentially different memory configurations and symbolic variables (e.g. i + 1 vs. i in the channel from operator 6 to 1).
[…]
$\mathrm{mem}_{\mathrm{llvm}}[\texttt{\%B}[\texttt{\%i}]\mapsto\texttt{\%A}[\texttt{\%i}]+1]=\mathrm{mem}_{\mathrm{df}}[\texttt{B}[\texttt{i}]\mapsto\texttt{A}[\texttt{i}]+1]$
%i + 1=i + 1%i + $=\texttt{i + 1}$ i + 1
If this check succeeds for all dataflow and LLVM cut points, we soundly conclude that we have obtained a bisimulation between the LLVM program and the dataflow program (under the deterministic canonical schedule).
Confluence Check
Once we have established a bisimulation between the LLVM function and the canonical schedule of the dataflow program, we check that the dataflow program is confluent: all possible schedules converge to the canonical schedule in the final state (or synchronize with it infinitely often if the program is non-terminating).
[…]
We attach a permission variable (representing a permission token that we do not know a priori) to each value in the channels of dataflow configurations in each symbolic branch. Figure 3 shows an example of this attachment of permission variables and permission constraints. At the starting state cut point 2, we attach free permission variables $p_{1},p_{2},p_{3},p_{4}$ to values in the configuration. After the first step in which operators 1 and 2 fire, these permission variables are consumed by the two carry operators and we generate fresh permission variables $p_{5},p_{6},p_{7}$ for the new output values. We require that the permission variables are used in a linear fashion: operators cannot duplicate or generate permission tokens, instead, they have to be obtained from the inputs. So for the first step in Figure 3, we add two constraints $p_{5}+p_{6}\leq p_{1}+p_{3}$ and $p_{7}\leq p_{2}+p_{4}$ to say that the output permissions have be contained in the input permissions. Furthermore, if the operator is a load or a store, such as in Figure 3 steps 4 and 6, we require that the suitable permission (read for load, and write for store) should be present in the input permissions (e.g. $\texttt{write}\leq p_{13}+p_{17}$ at step 6 in Figure 3).
Finally, when the execution is finished and has either hit a final state or another cut point, we have collected a list of constraints involving a set of permission variables. If the configuration is matched to a cut point, we add additional constraints to make sure that the assignment of tokens at the cut point is consistent. For example in Figure 3, for the last configuration, we would add these additional constraints: $p_{1}=p_{19}$, $p_{2}=p_{18}$, $p_{3}=p_{8}$, $p_{4}=p_{9}$. We then solve for the satisfiability of these constraints with respect to the permission algebra ($(\{0,\texttt{read},\texttt{write}\},+,\leq)$ we have defined above.
If there is a satisfying assignment of the permission variables with permission tokens, then intuitively, the operators in the dataflow program are consistent in memory accesses, and there will not be any data races. In Section 5, we show that this implies that any possible schedule will converge to the canonical schedule we use. For the example in Figure 3, a satisfying assignment is $p_{2}=p_{7}=p_{15}=p_{16}=p_{17}=p_{18}=\texttt{write}$ with all other variables set to 00. This assignment essentially passes the write permission across the loop of operators 2, 5, 7, 8, 9.
4. Bisimulation on the Canonical Schedule
Now, we detail the first half of our translation validation technique, which proves via cut-bisimulation that the LLVM program is equivalent to the dataflow program when restricted to a canonical schedule.
Given a dataflow program and LLVM program, their semantics are defined by transition systems $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$, respectively, where $\mathcal{I}_{\mathrm{df}}$ and $\mathcal{I}_{\mathrm{llvm}}$ are the sets of initial configurations and $\mathcal{C}_{\mathrm{df}}$ and $\mathcal{C}_{\mathrm{llvm}}$ are the sets of all configurations.
We use $\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})$ to denote the state of the memory in a dataflow configuration $c_{\mathrm{df}}$, and similarly $\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})$ for the memory in an LLVM configuration $c_{\mathrm{llvm}}$. We will omit the subscripts when it is clear from the context. For the dataflow semantics, we use $\to^{o}_{\mathrm{df}}$ to denote the (deterministic) firing of a particular operator $o$. For a transition relation $\to$, we use $\to^{*}$ for the reflexive-transitive closure of $\to$, and $\to^{+}$ for the transitive closure of $\to$. A configuration $c$ is terminating/final if there exists no configuration $c^{\prime}$ such that $c\to c^{\prime}$.
One important tool that we use is symbolic execution, which is similar to concrete execution but performed with symbolic expressions in the configuration. In the following, we establish some notation for symbolic execution. Let $\mathbf{x}$ be a list of variables $\mathbf{x}=(x_{1},\dots,x_{n})$. Let $V$ be the set of all bit-vectors of any lengths. We use $V(\mathbf{x})$ to denote the set of symbolic expressions constructed from constants in $V$, variables in $\mathbf{x}$, and any bit-vector operation we could perform in our semantics (e.g. 32-bit zero bit-vector $\texttt{0}_{32}$, or terms like $x_{1}+\texttt{1}_{32}$ are in $V(\mathbf{x})$). For $s\in\{\mathrm{df},\mathrm{llvm}\}$, we use $\mathcal{C}_{s}(\mathbf{x})$ to represent the set of symbolic configurations with free variables $\mathbf{x}$ in the semantics $s$. A symbolic configuration $c(\mathbf{x})\in\mathcal{C}_{s}(\mathbf{x})$ is similar to a configuration but over symbolic expressions $V(\mathbf{x})$ instead of only $V$ (for example, a symbolic dataflow configuration can have symbolic expressions such as $x_{1}+\texttt{1}_{32}$ in the channels; a symbolic LLVM configuration can have symbolic expressions stored in variables as opposed to only concrete values in $V$). If we have a substitution $\sigma:\{x_{1},\dots,x_{n}\}\to V(\mathbf{y})$ from $\mathbf{x}$ to symbolic expressions over some other list of variables $\mathbf{y}$, we denote $c(\sigma)\in\mathcal{C}_{s}(\mathbf{y})$ to be the result of the substitution. We also lift the transition relation $\to_{s}$ (for $s\in\{\mathrm{df},\mathrm{llvm}\}$) to symbolic configurations: we use $c(\mathbf{x})\to_{s}c^{\prime}(\mathbf{x})\mid\varphi(\mathbf{x})$ to denote the symbolic execution from $c(\mathbf{x})$ to $c^{\prime}(\mathbf{x})$ with path condition $\varphi(\mathbf{x})$.
The dataflow transition relation $\to_{\mathrm{df}}$ could be nondeterministic (a configuration can have multiple outgoing transitions), while $\to_{\mathrm{llvm}}$ is deterministic. However, for this section, we only care about a canonical schedule for the execution of operators. We denote the transition relation restricted to this canonical schedule by $\to_{\mathrm{df\mid canon}}\subseteq\to_{\mathrm{df}}$, and we focus on this deterministic transition for this section. The exact meaning of this canonical schedule will be explained in Section 4.2, but essentially, it executes dataflow operators in the same order their corresponding LLVM instructions would execute.
The equivalence property that we want to check is that for any pair of concrete initial states $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}$ with $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$,
  • $c_{\mathrm{df}}\to^{*}_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$ for a final state $c^{\prime}_{\mathrm{df}}$ iff $c_{\mathrm{llvm}}\to^{*}_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$ for a final state $c^{\prime}_{\mathrm{llvm}}$, in which case $\mathrm{mem}(c^{\prime}_{\mathrm{df}})=\mathrm{mem}(c^{\prime}_{\mathrm{llvm}})$.
  • $c_{\mathrm{df}}$ is non-terminating iff $c_{\mathrm{llvm}}$ is non-terminating, and there are infinitely many pairs of states in their traces with equal memory.
Since the dataflow and LLVM traces could both be unbounded, naive symbolic execution from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{llvm}})$ may not terminate. So we need to prove a coinductive invariant about the pairs of traces starting from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{llvm}})$, and such an invariant is a cut-bisimulation relation.
Definition 0 (Cut-bisimulation).
Let $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}$ be a binary relation. We say that $\mathcal{R}$ is a cut-bisimulation iff for any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$, if $c_{\mathrm{df}}\to_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$ for some $c^{\prime}_{\mathrm{df}}$, then there are $(c^{\prime\prime}_{\mathrm{df}},c^{\prime}_{\mathrm{llvm}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{df}}\to^{*}_{\mathrm{df\mid canon}}c^{\prime\prime}_{\mathrm{df}}$ and $c_{\mathrm{llvm}}\to^{+}_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$. Conversely, if $c_{\mathrm{llvm}}\to_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$ for some $c^{\prime}_{\mathrm{llvm}}$, then there are $(c^{\prime}_{\mathrm{df}},c^{\prime\prime}_{\mathrm{llvm}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{llvm}}\to^{*}_{\mathrm{llvm}}c^{\prime\prime}_{\mathrm{llvm}}$ and $c_{\mathrm{df}}\to^{+}_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$.
This is a weaker variant of the usual notion of bisimulation where we only allow one step before synchronizing instead of allowing one or more steps. Work by Kasampalis et al. (keq, ) provides more detail about this variant of bisimulation.
To check equivalence, we look for cut-bisimulations that are memory-synchronizing:
Definition 0 ().
A binary relation $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}$ is memory-synchronizing iff
  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\subset\mathcal{R}$, $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$.
  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}$ with $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$, $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$.
  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})$ with $c_{\mathrm{df}},c_{\mathrm{llvm}}$ final and $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$, $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$.
If there exists a cut-bisimulation $\mathcal{R}$ that is memory synchronizing, then we can conclude via a coinductive argument that the two programs are equivalent.
In the following subsections, we first describe how we represent a cut-bisimulation relation via cut points (Section 4.1), and how to automatically check that a proposed relation is actually a cut-bisimulation via symbolic execution (Section 4.2). Then in Section 4.3, we describe how the cut points are inferred using hints from the compiler.
4.1. Describing a Cut-Bisimulation via Cut Points
To finitely describe a proposed cut-bisimulation, we use a list of pairs of dataflow and LLVM cut points. A cut point is a pair $P=(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{llvm}}(\mathbf{y}))\mid\varphi(\mathbf{x},\mathbf{y})$ of symbolic dataflow and LLVM configurations constrained by the correspondence condition $\varphi(\mathbf{x},\mathbf{y})$. In our case, the condition $\varphi(\mathbf{x},\mathbf{y})$ is a conjunction of equations between variables in $\mathbf{x}$ and $\mathbf{y}$. In particular, we require $\mathrm{mem}(c_{\mathrm{df}}(\mathbf{x}))=\mathrm{mem}(c_{\mathrm{llvm}}(\mathbf{y}))$ to be in $\varphi(\mathbf{x},\mathbf{y})$ for any cut point. Semantically, each cut point describes the relation $\mathcal{R}(P):=\{(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{llvm}}(\mathbf{y}))\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}\mid\mathbf{x},\mathbf{y}\in V,\varphi(\mathbf{x},\mathbf{y})\}$. For a list of $n$ cut points
$\{P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\}_{i=1}^{n},$
the relation they describe is the union of the subrelation each describes: $\mathcal{R}(P^{1},\dots,P^{n}):=\bigcup^{n}_{i=1}\mathcal{R}(P^{i})$.
To ensure that the cut-bisimulation we propose satisfies the property in Definition 2, we also require that the first cut point must be exactly the pairs of initial configurations with equal memory. That is $\mathcal{R}(P^{1})=\{(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}\mid\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})=\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})\}$. Furthermore, we require that the last cut point $P^{n}$ must be the pairs of final configurations with equal memory, that is $\mathcal{R}(P^{n})=\{(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}\mid c_{\mathrm{df}},c_{\mathrm{llvm}}\text{ final},\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})=\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})\}$. Then $\mathcal{R}(P^{1},\dots,P^{n})$ satisfies Definition 2.
4.2. Checking Cut-Bisimulation
Now that we have a way to symbolically describe a relation $\mathcal{R}(P^{1},\dots,P^{n})$ with cut points satisfying Definition 2, we need to check that it is indeed a cut-bisimulation.
To achieve this, we perform symbolic execution from each cut point and try to reach another cut point. Suppose we are checking cut point $P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$. We perform symbolic execution from $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ and $c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})$ until all of their symbolic branches hit another cut point. For the LLVM side, we get a list of $k_{i}$ symbolic branches with path conditions denoted by $\psi$:
$c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})\to^{+}_{\mathrm{llvm}}c^{i,1}_{\mathrm{llvm}}(\mathbf{y^{i}})\mid\psi_{\mathrm{llvm},1}(\mathbf{y^{i}})$
$\dots$
$c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})\to^{+}_{\mathrm{llvm}}c^{i,k_{i}}_{\mathrm{llvm}}(\mathbf{y^{i}})\mid\psi_{\mathrm{llvm},k_{i}}(\mathbf{y^{i}})$
where each right hand side $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})$ for $l\in\{1,\dots,k_{i}\}$ is an instance $c^{j}_{\mathrm{llvm}}(\sigma)$ of some other LLVM cut point $c^{j}_{\mathrm{llvm}}(\mathbf{y^{j}})$ with substitution $\sigma:\mathbf{y^{j}}\to V(\mathbf{y^{i}})$. Furthermore, since our LLVM semantics is deterministic, the path conditions should be disjoint, i.e., for any $l_{1},l_{2}\in\{1,\dots,k_{i}\}$, if $l_{1}\neq l_{2}$, $\psi_{\mathrm{llvm},l_{1}}(\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l_{2}}(\mathbf{y^{i}})$ is unsatisfiable.
For the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, since the dataflow semantics is nondeterministic, we try to replicate the execution of LLVM branches on the dataflow side. For each LLVM branch $l\in\{1,\dots,k_{i}\}$, we can get a trace of LLVM instructions that are executed from $c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})$ to $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})$. Using hints from the compiler, we know that a subset of these LLVM instructions map to a list of dataflow operators $(o^{l}_{1},\dots,o^{l}_{n_{l}})$. For each LLVM branch, to produce the corresponding dataflow branch, we symbolically execute the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ on the schedule $(o^{l}_{1},\dots,o^{l}_{n_{l}})$ with the path condition $\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$:
$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{o^{l}_{1}*}_{\mathrm{df}}\dots\to^{o^{l}_{n_{l}}*}_{\mathrm{df}}c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}})\mid\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$
where each $\to^{o*}_{\mathrm{df}}$ denotes the dataflow transition by first firing the operator $o$, and then firing any available steer operator until none is fireable anymore. The canonical schedule $\to_{\mathrm{df\mid canon}}$ is defined by this order of execution of operators.
Therefore, by replicating the execution of the LLVM branches from cut point $i$, we get a list of dataflow branches:
(1) $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df\mid canon}}c^{i,1}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},1}(\mathbf{x^{i}})$
$\dots$
$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df\mid canon}}c^{i,k_{i}}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},k_{i}}(\mathbf{x^{i}})$
where each path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}}):=\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$. We check that the dataflow symbolic execution from cut point $i$ has covered all concrete cases, i.e., all possible cases of $\mathbf{x^{i}}$, by checking the unsatisfiability of $\neg\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{df},l}(\mathbf{x^{i}})$.
Finally, we check that each pair of dataflow and LLVM branches $(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}}))$ matches some target cut point $P^{j}=(c^{j}_{\mathrm{df}}(\mathbf{x^{j}}),c^{j}_{\mathrm{llvm}}(\mathbf{y^{j}}))\mid\varphi^{j}(\mathbf{x^{j}},\mathbf{y^{j}})$:
$\{(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})\}\subseteq\mathcal{R}(P^{j})$
or equivalently, there are substitutions $\sigma^{i,l}_{\mathrm{df}}:\mathbf{x^{j}}\to V(\mathbf{x^{i}}),\sigma^{i,l}_{\mathrm{llvm}}:\mathbf{y^{j}}\to V(\mathbf{y^{i}})$ such that the following conditions hold:
  • The dataflow branch is an instance of the target dataflow cut point: $c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}})=c^{j}_{\mathrm{df}}(\sigma^{i,l}_{\mathrm{df}})$.
  • The LLVM branch is an instance of the target LLVM cut point: $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})=c^{j}_{\mathrm{llvm}}(\sigma^{i,l}_{\mathrm{llvm}})$.
  • The correspondence condition at the target cut point is satisfied given the source cut point correspondence and the path conditions:
    $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})\rightarrow\varphi^{j}(\sigma^{i,l}_{\mathrm{df}},\sigma^{i,l}_{\mathrm{llvm}})$
If the check described in this section is passed for all cut points $i$ and all branches $1,\dots,k_{i}$, we can conclude that $\mathcal{R}(P^{1},\dots,P^{n})$ is indeed a cut-bisimulation.
4.3. Heuristics to Infer Cut Points
The final step is to actually infer the LLVM and dataflow cut points. In general, program equivalence for Turing-complete models like LLVM and dataflow programs is undecidable (rogers-87, ), so our best hope is to have good heuristics to infer the cut points for the specific input/output programs of the compiler.
The placement of cut points is important for the symbolic execution to terminate. It is in theory correct (for terminating dataflow and LLVM programs) to have exactly two cut points: one at the initial configurations, and one for all final configurations. But the symbolic execution will try to simulate all possible traces together, which are unbounded, and the check may not terminate. To ensure termination, on the LLVM side, we place a cut point at each of the following program points:
  • The entry of the function (i.e. the initial configuration);
  • The back edge of each loop header block; and
  • The exit program point.
For the dataflow program, we need to place the first cut point at the initial configuration of the dataflow program, and the last cut point at the final configuration. For loops, ideally, we want to place the cut points at the corresponding “program points” at loop headers. However, due to the lack of control-flow in the dataflow program, the form of the corresponding dataflow configurations at these “program points” is hard to infer statically. Instead of generating the dataflow cut points ahead of time, we dynamically infer them during symbolic execution of dataflow branches. For example, after mirroring the LLVM branch execution in a dataflow branch such as in Equation 1, if the dataflow cut point corresponding to the LLVM cut point has not been inferred yet, we generalize the symbolic configuration on the RHS of the dataflow branch Equation 1 to a dataflow cut point by replacing symbolic expressions with fresh variables. The correspondence condition is then inferred using compiler hints to match fresh variables in the new dataflow cut point to an LLVM variable defined by some LLVM instruction.
Note that we do not need these heuristics to be sound. As long as the cut points for the initial and final dataflow/LLVM configurations are generated correctly and the cut-bisimulation check in Section 4.2 succeeds, the soundness of the cut-bisimulation follows. While we do not prove these heuristics correct, they work well in practice and cover all test cases in our evaluation (Section 6).
5. Confluence Checking using Linear Permission Tokens
After checking cut-bisimilarity between the LLVM program and the dataflow program on the canonical schedule, we still need to prove that this canonical schedule is general. Indeed, since the dataflow program is asynchronous and nondeterministic, other choices of schedule may lead to data races and inconsistent final states.
Therefore, in addition to finding a bisimulation on the canonical schedule, we want to prove a confluence property on the dataflow program:
Definition 0 ().
We say $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is (ground) confluent iff for any initial configuration $c\in\mathcal{I}_{\mathrm{df}}$, if $c\to^{*}_{\mathrm{df}}c_{1}$ and $c\to^{*}_{\mathrm{df}}c_{2}$, then there is $c^{\prime}\in\mathcal{C}_{\mathrm{df}}$ with $c_{1}\to^{*}_{\mathrm{df}}c^{\prime}$ and $c_{2}\to^{*}_{\mathrm{df}}c^{\prime}$.
In this section, we present a technique using what we call linear permission tokens to ensure that the dataflow program is confluent. Linear permission tokens are, in essence, a dynamic variant of the fractional permissions studied in the concurrent separation logic (fractional-perm, ; concurrent-separation-logic, ). In our case, their linear usage is dynamically enforced and the exact placement of permission tokens is resolved after symbolic execution.
We associate each value present in a dataflow execution trace with a permission token in the form of $\texttt{write}\ l$, $\texttt{read}\ l$, or any formal sum of them, for any memory location $l$. These tokens are used by dataflow operators in a linear fashion: they cannot be replicated, and they are disjoint in the initial state. A write token can be split into $n$ many read tokens (for a predetermined parameter $n$), and inversely, $n$ of such read tokens can be combined to get back a write token. As shown later in this section, if there is a consistent assignment of these linear permission tokens, we can conclude that there are no data races possible in any schedule, and thus all final states have to be the same.
In Section 5.1, we prove that the technique works in a bounded setting for a concrete trace $\tau$: if $\tau$ has a consistent assignment of permission tokens, then for any other trace $\tau^{\prime}$ bounded by $\tau$ (meaning that for each operator, $\tau^{\prime}$ executes it for an equal or smaller number of times than $\tau$), $\tau^{\prime}$ will converge to the final configuration of $\tau$.
Then in Section 5.2, we explain how this bounded confluence check can be extended to a full confluence check for the entire dataflow program by using the cut points that we have placed for the dataflow program.
5.1. Confluence on Bounded Traces
In this section, we show that if we have a $n$-step trace $c_{1}\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}$ (where $o_{1},\dots,o_{n}$ are the operators fired in each step) with a consistent assignment of permission tokens to each value, then for any other trace $c_{1}\to^{o^{\prime}_{1}}_{\mathrm{df}}\dots\to^{o^{\prime}_{m}}_{\mathrm{df}}c_{m+1}$ (not necessarily with a consistent assignment of permission tokens) such that the multiset $\{o^{\prime}_{1},\dots,o^{\prime}_{m}\}\subseteq\{o_{1},\dots,o_{n}\}$, we have $c_{m+1}\to^{*}_{\mathrm{df}}c_{n+1}$.
Let us first describe the permission tokens and their assignment to values. The permission tokens are drawn from the permission algebra defined below.
Definition 0 (Permission Algebra).
Let $L$ be a set of memory locations and $k$ be a positive integer. The permission algebra is a partial algebraic structure $(P(L,k),0,+,\leq)$ where
  • $(P(L,k),0,+)$ is the term algebra of the formal sums $\texttt{read}\ l_{1}+\dots+\texttt{read}\ l_{n}$, with $l_{1},\dots,l_{n}\in L$, modulo the commutativity and associativity of $+$, and we require that for a location $l$, at most $k$ copies of $\texttt{read}\ l$ is present in the formal sum. We call $p+q$ the disjoint union of $p,q\in P(L,k)$. We also denote
    $\texttt{write}\ l:=\underbrace{\texttt{read}\ l+\dots+\texttt{read}\ l}_{k\text{ copies}}$
  • $P(L,k)$ is partially ordered by $\leq$ with $p\leq q$ (reads $q$ contains $p$) iff for any $l$, the number of occurrences of $\texttt{read}\ l$ in $p$ is less than or equal to that in $q$.
In Boyland’s terminology (fractional-perm, ), $P(L,k)$ is a finite token model of size $k$ of fractional permissions. For example, in $P(\{\texttt{A},\texttt{B}\},2)$, we have that
[…]
Definition 0 ().
We say that the permissions $p_{1},\dots,p_{n}\in P(L,k)$ are disjoint iff for any $i\neq j$, $p_{i}+p_{j}$ is defined.
Now let us define an assignment of permission tokens to values in a configuration. Let $E$ denote the set of channel names. For a configuration $c\in\mathcal{C}_{\mathrm{df}}$ and a channel $e\in E$, let $\mathrm{Channel}(c,e)$ denote the state of the channel $e$ in $c$, which is a string over $V$ representing the values in the channel $e$.
Definition 0 (Permission Augmentation).
Let $c\in\mathcal{C}_{\mathrm{df}}$ be a dataflow configuration. A permission augmentation of $c$ is a partial map $t:E\times\mathbb{N}\to P(L,k)$ such that for any $e\in E$, $t(e,n)$ is defined iff $n\in\{1,\dots,|\mathrm{Channel}(e)|\}$. We call $(c,t)$ a permission-augmented configuration.
Intuitively, $t(e,n)$ is the permission we attach to the $n$-th value in the channel $e$. In the following, we define when such a permission augmentation is consistent (i.e., the use of permissions is linear).
Definition 0 (Consistent Augmentation).
We call $(c,t)$ a consistent (permission-augmented) configuration if the permissions in the image of $t$ are disjoint.
Definition 0 (Consistent Transition).
Let $c\to^{o}_{\mathrm{df}}c^{\prime}$ be a transition. Let $t,t^{\prime}$ be two permission augmentations for $c,c^{\prime}$, respectively. Let $p_{1},\dots,p_{n}$ be the permissions in $t$ attached to input values of $o$ and $q_{1},\dots,q_{m}$ be the permissions in $t^{\prime}$ attached to output values of $o$. We say $(c,t)\to^{o}_{\mathrm{df}}(c^{\prime},t^{\prime})$ is a consistent (permission-augmented) transition iff:
  • The $t$ and $t^{\prime}$ have the same assignment of permissions except for the channels changed by the operator $o$.
  • (Linearity) $q_{1}+\dots+q_{m}\leq p_{1}+\dots+p_{n}$.
  • (Load Permission) If $o$ is a load operator on a memory location $l\in L$, then $\texttt{read}\ l\leq p_{1}+\dots+p_{n}$.
  • (Store Permission) If $o$ is a store operator on a memory location $l\in L$, then $\texttt{write}\ l\leq p_{1}+\dots+p_{n}$.
Definition 0 (Consistent Trace).
Let $c_{1}\to_{\mathrm{df}}\dots\to_{\mathrm{df}}c_{n}$ be a trace. Let $t_{1},\dots,t_{n}$ be permission augmentations to $c_{1},\dots,c_{n}$, respectively. We say that $(c_{1},t_{1})\to_{\mathrm{df}}\dots\to_{\mathrm{df}}(c_{n},t_{n})$ is a consistent (permission-augmented) trace iff:
  • $(c_{1},t_{1})$ is consistent
  • $(c_{i},t_{i})\to_{\mathrm{df}}(c_{i+1},t_{i+1})$ is consistent for $i\in\{1,\dots,n\}$.
The following lemma allows us to swap the order of execution of operators in two consistent steps.
Lemma 8 ().
Let $(c_{1},t_{1})\to^{o_{1}}_{\mathrm{df}}(c_{2},t_{2})\to^{o_{2}}_{\mathrm{df}}(c_{3},t_{3})$ be a consistent trace. Suppose $c_{1}\to^{o_{2}}_{\mathrm{df}}c^{\prime}_{2}$ for some configuration $c^{\prime}_{2}$. Then $c^{\prime}_{2}\to^{o_{1}}_{\mathrm{df}}c_{3}$ and there is a permission augmentation $t^{\prime}_{2}$ such that $(c_{1},t_{1})\to^{o_{2}}_{\mathrm{df}}(c^{\prime}_{2},t^{\prime}_{2})\to^{o_{1}}_{\mathrm{df}}(c_{3},t_{3})$ is consistent.
Proof Sketech.
If one of $o_{1}$ and $o_{2}$ is neither load nor store, or if both of them are loads, the result will always be consistent.
If $o_{1}$ and $o_{2}$ are both stores, and they are the same node, then it must be the case that $c^{\prime}_{2}=c_{2}$ and we can use $t^{\prime}_{2}=t_{2}$.
Otherwise, assume both $o_{1}$ and $o_{2}$ are load/store operators, and they are different nodes in the dataflow graph. Since they can both fire at $c_{1}$, and they are different nodes, their input values must be attached with disjoint permissions. Therefore, $o_{1}$ and $o_{2}$ have no dependencies, and they will give the same result no matter in which order we execute them. ∎
Inductively applying this lemma, we can swap the execution of an operator all the way to the beginning if it can fire and has not fired yet.
Lemma 9 ().
Let $(c_{1},t_{1})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}(c_{n+1},t_{n+1})$ be a consistent trace. Suppose $o_{n}$ is not in the set $\{o_{1},\dots,o_{n-1}\}$ and $c_{1}\to^{o_{n}}_{\mathrm{df}}c^{\prime}_{2}$ for some $c^{\prime}_{2}$, then $c_{1}\to^{o_{n}}_{\mathrm{df}}c^{\prime}_{2}\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n-1}}_{\mathrm{df}}c_{k}$ for some $c^{\prime}_{2}$ and there is a consistent augmentation for this trace.
Applying the lemma above inductively, we have the following.
Theorem 10 (Bounded Confluence for Consistent Traces).
Let $(c_{1},t_{1})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}(c_{n+1},t_{n+1})$ be a consistent trace. Let $c_{1}\to^{o^{\prime}_{1}}_{\mathrm{df}}\dots\to^{o^{\prime}_{m}}_{\mathrm{df}}c_{m+1}$ be another trace. If we have the multiset inclusion $\{o^{\prime}_{1},\dots,o^{\prime}_{m}\}\subseteq\{o_{1},\dots,o_{n}\}$, then $c_{m+1}\to^{*}_{\mathrm{df}}c_{n+1}$.
Practically, with this bounded confluence theorem, we can do bounded model checking for the confluence property: we perform symbolic execution to obtain a trace of symbolic configurations $c_{1}(\mathbf{x})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}(\mathbf{x})$, and then we can solve for a consistent augmentation $(t_{1},\dots,t_{n+1})$ for this trace using an SMT (z3, ) solver. If there is a solution, then it follows from Theorem 10 that any other trace (firing operators only in the multiset $\{o_{1},\dots,o_{n}\}$, counting multiplicity) would still converge to $c_{n+1}(\mathbf{x})$ in the end.
5.2. Full Confluence Checking using Cut Points
To extend the bounded confluence results in Section 5.1 to checking confluence for the entire dataflow program with unbounded traces, we need to finitely describe and check a consistent pattern of permission augmentations for all traces. For this purpose, we use the cut points selected in Section 4.
Let $P^{1}_{\mathrm{df}}(\mathbf{x^{1}}),\dots,P^{n}_{\mathrm{df}}(\mathbf{x^{n}})$ be a list of cut points. During the bisimulation check in Section 4.2, we have verified that starting from any cut point $P^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, any branch of symbolic execution reaches an instance of another cut point $P^{j}_{\mathrm{df}}(\mathbf{x^{j}})$. A consequence of this check is that all the symbolic branches from all cut points “cover” any complete trace (traces from an initial configuration to a final configuration). In other words, any complete concrete trace from the initial configuration is the concatenation of instances of symbolic branches.
Thus, if we are able to find a consistent permission augmentation for the trace of each symbolic branch during the bisimulation check, then any concrete trace in $\to_{\mathrm{df\mid canon}}$ would have a consistent permission augmentation, which implies that any other schedule will converge to $\to_{\mathrm{df\mid canon}}$, and the confluence of the entire dataflow program follows.
Theorem 11 ().
If all symbolic dataflow branches produced in Section 4.2 have consistent permission augmentations, then $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is confluent.
Figure 5. Evaluation results on a benchmark of 20 programs. From left to right, the columns are names of the test cases (Name), lines of code in the original C program (LOC), numbers of dataflow operators (#OP), numbers of permission constraints (#PC), time spent for bisimulation check in seconds (Bisim.), time spent for confluence check in seconds (Conf.). All test cases are written originally in C and then compiled to LLVM via Clang (clang, ). The first 10 programs are benchmark programs from the RipTide work. The nn_* tests implement neural network inference. The last four are selected from a follow-up work to RipTide called Pipestitch (pipestitch, ). In the case of sort and SpSlice, the bisimulation check fails, because they contain instances of the same compiler bug. In all other cases, the bisimulation and confluence checks succeed.
Name LOC #OP #PC Bisim. (sec) Conf. (sec)
bfs 48 48 3634 0.45 7.79
dconv 24 65 8134 0.96 11.87
dfs 45 49 3762 0.46 8.51
dmm 16 56 4220 0.71 6.31
dmv 10 31 1234 0.22 1.69
fft 29 70 2690 0.45 6.04
sconv 22 56 4272 0.71 10.89
smm 22 49 3967 0.53 12.08
smv 12 25 1103 0.19 2.36
sort 23 57 6554 $\times$ 0.64 3.62
nn_conv 43 61 12154 0.78 15.65
nn_fc 24 36 2632 0.31 3.43
nn_norm 15 19 316 0.07 0.17
nn_pool 33 52 8165 0.74 7.11
nn_relu 16 19 844 0.11 0.44
nn_vadd 9 15 244 0.06 0.15
Dither 21 31 2553 0.32 2.36
SpMSpMd 42 60 7693 0.93 28.45
SpMSpVd 48 40 3182 0.4 10.22
SpSlice 33 31 2616 $\times$ 0.34 7.4
Evaluation results.
Figure 5. Evaluation results on a benchmark of 20 programs. From left to right, the columns are names of the test cases (Name), lines of code in the original C program (LOC), numbers of dataflow operators (#OP), numbers of permission constraints (#PC), time spent for bisimulation check in seconds (Bisim.), time spent for confluence check in seconds (Conf.). All test cases are written originally in C and then compiled to LLVM via Clang (clang, ). The first 10 programs are benchmark programs from the RipTide work. The nn_* tests implement neural network inference. The last four are selected from a follow-up work to RipTide called Pipestitch (pipestitch, ). In the case of sort and SpSlice, the bisimulation check fails, because they contain instances of the same compiler bug. In all other cases, the bisimulation and confluence checks succeed.
6. Implementation and Evaluation
We have implemented our translation validation technique in the tool WaveCert  [^2] Available on GitHub at https://github.com/secure-foundations/riptide-verification.. We first implemented symbolic executors for (a subset of) LLVM and dataflow programs in Python (z3, ). Based on these, we implemented the bisimulation check in Section 4 and the confluence check in Section 5.
WaveCert uses the Z3 SMT solver (z3, ) to discharge the verification conditions it generates. In the bisimulation check, WaveCert encodes the validity of the correspondence conditions at each cut point into Z3 queries. In the confluence check, WaveCert produces a list of permission constraints and checks if these constraints are satisfiable with respect to the finite permission algebra (Definition 2) by encoding each permission variable as a set of Boolean variables and the list of permission constraints as a SAT query.
In order to evaluate WaveCert’s capability to certify compilation to real dataflow programs, we have applied WaveCert to a benchmark of 20 programs consisting of the benchmark programs in RipTide (riptide, ) and Pipestitch (pipestitch, ), as well as some programs implementing neural network inference. Figure 5 shows some statistics about the benchmark programs and how WaveCert performs. All tests are performed on a laptop with an M1 Max processor and 64 GiB of RAM.
In all of the benchmark programs, the bisimulation check either succeeds or finds an actual bug (such as in sort and SpSlice, which we discuss further in Section 6.1). This shows that our heuristics for inferring dataflow cut points works well on the canonical schedule. Our bisimulation check generates verification conditions in quantifier-free theories of arrays and integers without non-linear arithmetic, so all bisimulation checks finish quickly within a second.
The confluence check succeeds in all examples. For sort, we manually fixed the incorrect dataflow program produced by the compiler, after which the confluence check passes.
Figure 6. Examples of two compiler bugs.
Examples of two compiler bugs.
Figure 6. Examples of two compiler bugs.
6.1. Compiler Bugs
During the testing of WaveCert, we found 7 bugs in the RipTide compiler. Three of these bugs were discovered on the test cases sort, SpSplice, and their variations; and the rest four bugs were found on smaller test cases that we manually constructed. A developer of the RipTide compiler has confirmed these issues.
In the following, we discuss two examples of these bugs to show that the bisimulation check and the confluence check are both effective in catching compilation bugs.
Incorrect signed extension
WaveCert found a compilation bug involving integer signed vs. unsigned extension in the sort test case. A simplified LLVM program used to reproduce this issue is shown in the top left of Figure 6, and the output dataflow program is shown in the top right of Figure 6. The issue itself is straightforward: in the LLVM code, a 1-bit xor is done with a 1-bit constant 1, and then the result is zero-extended to a 32-bit integer. On the dataflow side, the dataflow compiler tries to convert all operations into 32-bits, the word width supported on the RipTide architecture. However, during the process, the compiler assumes that all constants should be sign-extended. As a result, it extends the 1-bit 1 to a 32-bit -1, which is incorrect in this case. Our bisimulation check catches this bug when verifying the correspondence condition at the final cut point, where the memory configurations do not match.
Incorrect memory ordering with potential data races
Our confluence check catches a compilation issue when compiling from the C program shown in the bottom left of Figure 6 to the dataflow program in the bottom right of Figure 6. In the C program, we have a nested loop. In each iteration of the outer loop, we first set A[i] = i, and then in the inner loop that follows, we copy A[i] into A[i + 1], …, A[n - 1]. In the output dataflow program, the compiler allows the load and store operators in the inner loop to pipeline using the invariant operator (marked I) to repeatedly send the signal to enable load. However, this causes a data race, because in the inner loop, a load in the second iteration accesses the same memory location as the store in the first iteration.
Our confluence check fails as the list of permission constraints we collected are unsatisfiable. WaveCert also outputs an unsat core of 57 constraints, a subset of the 420 constraints generated, that causes the unsatisfiability. In particular, the unsat core includes two ordering constraints: read  A is contained in the disjoint union of input tokens of the inner loop load and write  A is contained in the disjoint union of input tokens of the inner loop store. This indicates that the main conflict stems from the inner loop load and store.
7. Limitations and Future Work
Our translation validation technique still has some limitations, and we plan to address them in future work.
Optimized dataflow graphs
Our technique may fail on more optimized programs. The RipTide compiler employs some optimizations currently not supported by WaveCert:
  • Deduplication of operators, which combines arithmetic operators with the same inputs.
  • Streamification, which uses a single stream operator to generate a loop induction variable (as opposed to using, e.g., carry, comparison, and addition operators).
  • Array dependency analysis, which removes unnecessary orderings between loads and stores across iterations that access different indices of an array.
The first two optimizations are not supported by our bisimulation check because we require a one-to-one correspondence between LLVM instructions and dataflow operators for the canonical schedule. In future work, we can relax this requirement and allow multiple LLVM instructions to map to one dataflow operator, and we then only fire the operator if all of the corresponding LLVM instructions have fired.
The third optimization may lead to a dataflow program failing our confluence check. To solve this, we need finer granularity in permission tokens. Currently, the read and write permissions are for the entire memory region, such as an array A. To allow memory operations on A[i] and A[i + 1] to run in parallel, where i is the loop induction variable, we need to re-formulate the permission algebra to allow slices of a permission token on a specific range of indices. This also requires a form of “dependent” permission tokens that can depend on the values in the channels.
Liveness properties
Our semantics of the dataflow program uses unbounded FIFO queues to model dataflow channels. In this model, we do not need to consider liveness issues such as deadlocks, because an operator can always fire once the input values are ready. However, this is not true for the actual hardware, which has a limit on the queue size.
For future work, we plan to explore model checking (model-checking, ) techniques to check for deadlocks.
Lack of architecture-specific details
This work targets the compilation from imperative programs to an abstract version of dataflow programs. While our technique can ensure the correctness of this pass, we do not verify later passes, such as the procedure to optimally configure the dataflow program on the actual hardware.
8. Related Work
As a compiler verification technique, the translation validation approach taken in our work is a less labor-intensive alternative to full compiler verification, such as CompCert (compcert, ). It was first proposed by Pnueli et al. (pnueli-tv, ), and there are numerous following works with novel translation validation or program equivalence checking techniques for various languages and compilers (necula-tv, ; kundu-tv, ; data-driven-eq, ; tate-eq-saturation, ). For LLVM-related works on translation validation, we have Alive (alive, ; alive2, ), which performs bounded translation validation for internal optimizations in LLVM. LLVM-MD (llvm-md, ; denotational-tv, ) is another translation validation tool for LLVM optimizations. LLVM-MD translates the LLVM code into an intermediate representation called synchronous value graphs (SVG) (denotational-tv, ), which is a similar representation to a dataflow program. The difference is that the semantics of SVG is synchronous, and they also trust the compilation from LLVM to SVG, whereas in our case, such translation is exactly what we are trying to validate. Our formulation of cut-bisimulation in our bisimulation check is an adaptation from the work on Keq (keq, ), which aims to be a more general program equivalence checker parametric in the operational semantics of input/output languages.
Our use of linear permission tokens is a dynamic variant of the fractional permissions proposed by Boyland (fractional-perm, ) and used in some variations of concurrent separation logic (concurrent-separation-logic, ). In contrast to separation logic frameworks such as Iris (iris, ), which are embedded in interactive proof assistants, our permission tokens are synthesized and checked fully automatically.
There are also verification tools for classical dataflow languages such as Lustre (lustre, ; lustre-verified-compiler, ; lustre-smt-verification, ) and Esterel (esterel-verification, ). These languages have a synchronous semantics and do not interact with globally accessible memory, whereas dataflow programs emitted by CGRA compilers are highly asynchronous and make pervasive use of the global heap.
In the context of term rewriting systems (term-rewriting, ), the standard approach for confluence checking is the use of critical pairs (knuth-critical-pairs, ; duran-confluence, ), which proves ground confluence in the case when the rewriting system is terminating. This algorithm is implemented in various systems, such as the Church-Rosser checker in Maude (maude, ). However, this technique is difficult to apply in our case, since the semantics of a dataflow program is neither terminating nor, in general, ground confluent.
9. Conclusion
In this work, we develop a translation validation technique for asynchronous dataflow programs with an application to the RipTide CGRA architecture. Our technique first checks the bisimulation of the input imperative program and the output dataflow program on a canonical schedule, and then we generalize the results to all possible schedules via linear permission tokens. Our work is an initial step towards integrating high assurance into new architectures from the start, rather than retrofitting them decades later.

### v2

@@
-WaveCert: Formal Compiler Validation for Asynchronous Dataflow Programs
+WaveCert: Translation Validation for Asynchronous Dataflow Programs via Dynamic Fractional Permissions
 Abstract.
-Coarse-grained reconfigurable arrays (CGRAs), or dataflow architectures in general, have gained attention in recent years due to their promising power efficiency compared to traditional von Neumann architectures.
+Coarse-grained reconfigurable arrays (CGRAs) have gained attention in recent years due to their promising power efficiency compared to traditional von Neumann architectures.
 To program these architectures using ordinary languages such as C, a dataflow compiler must transform the original sequential, imperative program into an equivalent dataflow graph, composed of dataflow operators running in parallel.
@@
 We apply this method to a state-of-the-art dataflow compiler targeting the RipTide CGRA architecture.
-Our tool uncovers 7 compiler bugs where the compiler outputs incorrect dataflow graphs, including a data race that is otherwise hard to discover via testing.
+Our tool uncovers 8 compiler bugs where the compiler outputs incorrect dataflow graphs, including a data race that is otherwise hard to discover via testing.
 After repairing these bugs, our tool verifies the correct compilation of all programs in the RipTide benchmark suite.
+copyright: none
 1. Introduction
 Even as Moore’s Law ends, many applications still demand better power efficiency, e.g., for high-performance machine learning or for ultra-low-power environmental monitoring.
-A flood of new hardware architectures have emerged to meet this demand, but one attractive approach that retains general-purpose programmability is a class of dataflow architectures called coarse-grained reconfigurable arrays (CGRAs) (cgra-survey, ).
+A flood of new hardware architectures has emerged to meet this demand, but one attractive approach that retains general-purpose programmability is a class of dataflow architectures called coarse-grained reconfigurable arrays (CGRAs) (Liu et al., 2019).
 In a CGRA architecture, an array of processing elements (PEs) is physically laid out on the chip.
 Each PE can be configured to function as a high-level operator that performs an arithmetic operation, a memory operation, or a control-flow operation.
-These PEs are interconnected via an on-chip network, reducing communication and data movement costs compared to traditional von Neumann architectures, thus making a CGRA more power efficient.
+These PEs communicate via an on-chip network, reducing data movement costs compared to traditional von Neumann architectures, thus making a CGRA more power efficient.
 Abstractly, CGRAs run dataflow programs, or networks of operators that execute independently and asynchronously.
 While CGRAs are able to run dataflow programs efficiently due to their inherent parallelism, this parallelism also introduces new, subtle correctness issues due to the potential for data races.
-As a result, dataflow programs are not written directly, but are instead compiled from ordinary, imperative languages.
+As a result, dataflow programs are not written directly, but are instead compiled from sequential, imperative languages.
 However, the gap between imperative and dataflow programming means that these compilers are complex; to maximize performance, they must use a variety of program analyses to convert imperative programs into equivalent distributed ones.
@@
 Since CGRAs cannot execute as intended without correct dataflow programs, we begin this effort by verifying the translation from imperative programs to dataflow programs.
-In this work, we propose using translation validation (pnueli-tv, ) to certify the results of dataflow compilers.
+In this work, we propose using translation validation (Pnueli et al., 1998) to certify the results of dataflow compilers.
 In our case, translation validation amounts to checking that the output dataflow program from the compiler has the same behaviors as the input imperative program.
@@
 Additionally, in contrast to full compiler verification, translation validation allows the use of unverified compilation toolchains, which is crucial for the fast-developing area of CGRA architectures.
-Our main challenge for leveraging translation validation is managing the asynchronicity of dataflow programs.
-Most translation validation works (pnueli-tv, ; necula-tv, ; kundu-tv, ; data-driven-eq, ) prove equivalence between input and output programs via (bi)simulation, or relational invariants between the states of the two programs.
-Since dataflow programs have exponentially more states than imperative programs (arising from asynchronous scheduling decisions), constructing this bisimulation relation directly would be quite tricky.
+Our main challenge for leveraging translation validation is managing the asynchrony of dataflow programs.
+Most translation validation works (Pnueli et al., 1998; Necula, 2000; Kundu et al., 2009; Sharma et al., 2013) prove equivalence between input and output programs via simulation, or relational invariants between the states of the two programs.
+Since dataflow programs have exponentially more states than imperative programs (arising from asynchronous scheduling decisions), constructing this simulation relation directly would be quite tricky.
 Instead, we utilize a two-phase approach.
-First, we use symbolic execution to prove that there is a bisimulation between the input imperative program and the output dataflow program on a canonical schedule.
+First, we use symbolic execution to prove that there is a simulation between the input imperative program and the output dataflow program on a canonical schedule.
 In the canonical schedule, the dataflow operators are scheduled to fire in a similar order to their counterparts in the original imperative program.
-Next, we prove that any possible schedule of the dataflow program must converge to the same final state as the canonical schedule; i.e., it is confluent (term-rewriting, ).
+Next, we prove that any possible schedule of the dataflow program must converge to the same final state as the canonical schedule; i.e., it is confluent (Baader and Nipkow, 1998).
 For our setting, proving confluence requires showing that the dataflow program does not contain any data races.
-Taking inspiration from separation logic (separation-logic, ), we augment our symbolic execution with linear permission tokens to track ownership of memory locations.
-Taken together with the initial bisimulation phase, we can prove that the imperative program is equivalent to the dataflow program on all possible inputs and schedules.
-We have implemented this technique in WaveCert  [^1] Available on GitHub at https://github.com/secure-foundations/riptide-verification., a translation validation system targeting the state-of-the-art CGRA architecture RipTide (riptide, ).
-RipTide features a dataflow compiler from LLVM (llvm, ) programs to dataflow programs.
-Using our tool, we found 7 compiler bugs where the RipTide dataflow compiler generated incorrect dataflow programs.
-All of these bugs were confirmed by the developer of the RipTide compiler.
+Inspired by fractional permissions (Boyland, 2013), we augment our symbolic execution with linear permission tokens to track ownership of memory locations.
+In contrast with interactive program logics (e.g., concurrent separation logic (Brookes, 2006)) that typically require manual annotations, we automatically compute dynamic permissions that may flow in arbitrary (sound) ways.
+Putting together both phases, we prove that the imperative program is equivalent to the dataflow program on all possible inputs and schedules.
+Furthermore, equivalence with the sequential program also implies liveness and deadlock freedom for the dataflow program, as it enforces that the dataflow program can make progress whenever the sequential program can.
+We have implemented this technique in WaveCert, a translation validation system targeting the state-of-the-art CGRA architecture RipTide (Gobieski et al., 2023), which operates via a compiler from LLVM (Lattner and Adve, 2004) programs to dataflow programs.
+Using our tool, we found 8 compiler bugs where the RipTide dataflow compiler generated incorrect dataflow programs.
+All of these bugs were confirmed by the developers of the RipTide compiler.
 One of these bugs allows the compiler to emit a dataflow program with data races, which are hard to discover via testing and costly to fix after deployment.
 To summarize, our contributions in this work are:
-  • A novel two-phase translation validation technique for certifying compilation results from imperative programs to dataflow programs;
-  • An implementation of our technique, WaveCert, which targets the RipTide dataflow compiler; and
-  • An evaluation of WaveCert on the RipTide dataflow compiler, which revealed 7 compilation bugs. Each successful verification result takes less than 30 seconds, with most taking around 10 seconds.
+  • A novel two-phase translation validation technique to prove that the output dataflow program has equivalent behavior as the input imperative program, capturing both correctness and liveness properties.
+  • An implementation of our technique, WaveCert  [^1] Source code is available in an GitHub repository (Lin et al., 2024), targeting the RipTide dataflow compiler.
+  • An evaluation of WaveCert on the RipTide dataflow compiler, which reveals 8 compilation bugs. Most verification result takes around 10 seconds with a maximum of about 30 seconds.
+[…]
 Examples of operators.
 […]
-2. Preliminaries: CGRAs and RipTide
-Dennis et al.
-(dennis-dataflow, ) introduced early designs for dataflow programming and architectures in the mid 1970s to allow programs to achieve a higher degree of concurrency.
-In the intervening years, the community has explored a wide variety of dataflow architectures (dataflow-architectures, ; yazdanpanah-dataflow, ).
-The desire for low-power, general-purpose programmability has recently led to a line of work on coarse-grained reconfigurable arrays (CGRAs) (cgra-survey, ).
-To utilize the efficiency and parallelism in CGRAs, the predominant general-purpose way to program CGRAs is to compile from a sequential, imperative program into a dataflow graph via a dataflow compiler.
-This is used in CGRA projects such as PipeRench (piperench, ), WaveScalar (wavescalar, ), and more recently, RipTide (riptide, ) and Pipestitch (pipestitch, ).
-In this work, we focus on the approach taken by RipTide (riptide, ), which has a dataflow compiler from arbitrary LLVM functions to dataflow programs.
-For an informal description of the semantics of the LLVM intermediate representation (IR), we refer the reader to the official LLVM IR manual (llvm-ir, ).
-In this work, we use the subset of LLVM IR that is supported by the RipTide compiler, including basic control-flow and integer arithmetic.
-The dataflow program output by the RipTide compiler is a Turing-complete representation of programs.
+2. Preliminaries: Dataflow Programs, CGRAs, and RipTide
+Dataflow programming (Johnston et al., 2004) is an old idea dating back to the 1960s and 1970s.
+Early foundational works include the dataflow architecture by Dennis et al.
+(Dennis and Misunas, 1974), as well as theoretical analysis of models of dataflow programs such as computation graphs (Karp and Miller, 1966) and Kahn process networks (Gilles, 1974).
+The main motivation in these early works is to achieve a greater degree of instruction-level parallelism (Ackerman, 1982).
+In the intervening years, the community has explored a wide variety of dataflow architectures (Arvind and Culler, 1986; Yazdanpanah et al., 2014).
+Among these dataflow architectures, coarse-grained reconfigurable arrays (CGRAs) (Liu et al., 2019) stand out due to their promising power efficiency: recent work (Gobieski et al., 2021, 2023) is able to support general programmability, while reducing the energy cost by orders of magnitude compared to a von Neumann core.
+To utilize the efficiency and parallelism in CGRAs, the predominant general-purpose way to program CGRAs is to compile a sequential, imperative program into a dataflow graph via a dataflow compiler, as seen in CGRA projects such as PipeRench (Goldstein et al., 2000), WaveScalar (Swanson et al., 2003), and more recently, RipTide (Gobieski et al., 2023) and Pipestitch (Serafin et al., 2023).
+In this work, we focus on the approach taken by RipTide (Gobieski et al., 2023), which has a dataflow compiler from arbitrary LLVM functions to dataflow programs.
+We briefly introduce LLVM and dataflow programs below.
+LLVM IR
+LLVM (Lattner and Adve, 2004) is a compiler framework and features an intermediate representation called LLVM IR.
+Compiler writers can translate a higher-level language such as C or Rust into LLVM IR, and then the LLVM framework can perform optimizations on LLVM IR and produce machine code in various target architectures.
+For an informal description of the semantics of the LLVM IR, we refer the reader to the official LLVM IR manual (LLVM, 2024b).
+In this work, we use the subset of LLVM IR supported by the RipTide compiler, including basic control-flow, integer arithmetic, and memory operations.
+Function calls and floating-point operations are not supported.
+Dataflow programs
+A dataflow program is a Turing-complete representation of programs.
 It is represented as a dataflow graph in which the nodes are called operators and edges are called channels.
-Semantically, channels are unbounded FIFO queues used for communication between operators.
-Operators can be thought of as processes running in parallel with a local state.
-Repeatedly, each operator: waits to dequeue (a subset of) input channels to get input values; performs a local computation, optionally changing the local state or global memory; and pushes output values to (a subset of) output channels.
+Semantically, operators can be thought of as stateful processes running in parallel, which communicate through channels, or FIFO queues of messages.
+Repeatedly, each operator: waits to dequeue (a subset of) input channels to get input values; performs a local computation, optionally changing its local state or global memory; and pushes output values to (a subset of) output channels.
 When an iteration of this loop is done, we say that the operator has executed or fired.
-We model the execution of dataflow operators using a fully asynchronous semantics.
-Different CGRAs would use different scheduling strategies (cgra-survey, ), so we conservatively use asynchronicity to cover all possible schedules.
-In Figure 1, we show the five most important types of operators.
-Reading from left to right, we first have the control flow operators, carry (C) and steer (T).
+Since different CGRAs have different strategies for scheduling operators (Liu et al., 2019), we conservatively use asynchrony to model all possible schedules.
+In Figure 1, we show the five most important types of operators in RipTide (Gobieski et al., 2023) [^2] WaveCert supports all RipTide operators, except for the Stream operator, as described in Section 7..
+In Figure 1, reading from left to right, we first have the control flow operators, carry (C) and steer (T).
 The carry operator is used to support loop variables.
@@
 In the loop state, it waits for values from B and D (decider); if D is true, it outputs O = B; otherwise it discards B and transitions to the initial state.
-The steer operator is used for conditional execution. It waits for values from A and D. If D is true, it outputs O = A; otherwise, it discards the value, and outputs nothing.
+The steer operator is used for conditional execution. It waits for values from A and D. If D is true, it outputs O = A; otherwise, it discards the value and outputs nothing.
 To utilize the global memory, we have the load (LD) and store (ST) operators.
@@
 To compile an LLVM program to a dataflow program, the RipTide compiler first maps each LLVM instruction to its corresponding dataflow operator, and then it enforces the expected control-flow and memory ordering semantics of the original program.
-Using a control dependency analysis, it inserts steer operators to selectively enable/disable operators depending on branch/loop conditions.
-Using a memory ordering analysis, it inserts dataflow dependencies between load/store operators to prevent data races.
+Using a control dependency analysis, it inserts steer operators to selectively enable/disable operators depending on branch/loop conditions; Using a memory ordering analysis, it inserts dataflow dependencies between load/store operators to prevent data races.
 These analyses are quite complex and use various optimizations to allow more parallelism than the original LLVM program.
 As shown in our evaluation in Section 6, this process can easily have bugs.
-Hence, in our work, we take the translation validation approach to certify the correctness of dataflow compilation.
-Figure 2. An example of input/output programs from the RipTide dataflow compiler.
+Hence, in our work, we use translation validation to certify the correctness of dataflow compilation.
+3. Overview and an Example
+WaveCert performs translation validation on the RipTide dataflow compiler, which compiles LLVM programs to dataflow programs.
+Given the input program and the compiled output program (along with some hints generated by the compiler), WaveCert performs two checks: first, a simulation check, which verifies that the LLVM program is equivalent to the dataflow program on a restricted, canonical schedule of dataflow operators; and second, a confluence check, which proves that the choice of the canonical schedule for the dataflow program is general and all other schedules reach the same final state as the canonical schedule.
 1define void @test(i32* %A, i32* %B, i32 %len) {
@@
 3 br label %header
-4
-5header:
+4header:
 5 %i = phi i32 [ 0, %entry ], [ %i_inc, %body ]
@@
 7 br i1 %cond, label %body, label %end
-9
-10body:
+8body:
 9 %idx1 = getelementptr i32, i32* %A, i32 %i
@@
 15 br label %header
+16end:
+17 ret void
 18}
-19end:
-20  ret void
-21}
 (a) Input LLVM function.
 (b) Output dataflow program.
+Figure 2.
+An example of input/output programs from the RipTide dataflow compiler.
+Most dataflow operators correspond to LLVM instructions: the comparison operator corresponds to line 6, the add operators correspond to lines 11 and 14, while the load and store operators correspond to lines 10 and 13, respectively.
+[…]
 An example.
-Figure 2. An example of input/output programs from the RipTide dataflow compiler.
-Figure 3.
-A trace of execution from the dataflow cut point 2 following the canonical schedule of operators: 1, 2, 4, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2).
-In addition, steer operators 3 and 5 are also fired in step 3 immediately after they are ready.
-At the end of the trace, the configuration matches cut point 2 again.
-For each configuration, we show the state of the channels (solid lines for channels with one value, and dotted lines for empty channels).
-The lightning symbols indicate which operators are fired in the previous step.
-Before the firing of an operator, the input channels of the operator should be non-empty (i.e., marked with a solid line).
-After firing, the input channels are emptied, and the output channels will become non-empty.
-The exact symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked.
-At the top of each configuration, we also indicate the list of permission constraints we add at each step.
-Cut point 2 trace.
-Figure 3.
-A trace of execution from the dataflow cut point 2 following the canonical schedule of operators: 1, 2, 4, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2).
-In addition, steer operators 3 and 5 are also fired in step 3 immediately after they are ready.
-At the end of the trace, the configuration matches cut point 2 again.
-For each configuration, we show the state of the channels (solid lines for channels with one value, and dotted lines for empty channels).
-The lightning symbols indicate which operators are fired in the previous step.
-Before the firing of an operator, the input channels of the operator should be non-empty (i.e., marked with a solid line).
-After firing, the input channels are emptied, and the output channels will become non-empty.
-The exact symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked.
-At the top of each configuration, we also indicate the list of permission constraints we add at each step.
-[…]
-Index LLVM Dataflow Correspondence
-1 %entry Initial config
-$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
-2
-%body
-$\to$ %header
-See Figure 3
-$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
-[…]
-$\bot$ %end Final config
-$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
-Cut points for the example.
-[…]
-3. Overview and an Example
-WaveCert performs translation validation on the RipTide dataflow compiler, which compiles LLVM programs to dataflow programs.
-Given the input program and the compiled output program (along with some hints generated by the compiler), WaveCert performs two checks: first, a bisimulation check, which verifies that the LLVM program is equivalent to the dataflow program on a restricted, canonical schedule of dataflow operators; and second, a confluence check, which proves that the choice of schedule for the dataflow program does not affect the final result.
-If a pair of input/output programs passes both checks, the two programs are equivalent.
-If the bisimulation check fails, there might be a compilation bug that produced an incorrect dataflow.
-If the bisimulation check succeeds but the confluence check fails, there might be a data race in the output dataflow program.
+If a pair of input/output programs passes both checks, the two programs are equivalent (in the sense of Definition 2).
+Equivalence implies both safety — the dataflow program returns the correct values — and liveness — the dataflow program always terminates correctly, and does not contain communication-related deadlocks.
 Example
@@
 The loop increments variable %i from 0 to %len, and at each iteration, updates %B[%i] = %A[%i] + 1.
-The LLVM function @test gets compiled to the dataflow program in Figure 1(b), in which most operators correspond to an LLVM instruction:
-  • The comparison operator 4 corresponds to line 7;
-  • The load operator 7 corresponds to line 12;
-  • The add operator 8 corresponds to line 13;
-  • The store operator 9 corresponds to line 15; and
-  • The add operator 6 corresponds to line 16.
-The compiler also inserts additional operators to faithfully implement the sequential semantics of the LLVM program.
+The LLVM function @test is compiled to the dataflow program in Figure 2(b).
+In addition to compute and memory operators, the compiler also inserts operators to faithfully implement the sequential semantics of the LLVM program.
 The steer operators marked with T enforce that the operators corresponding to the loop body only execute when the loop condition is true.
-The carry operator 1 is used for the loop induction variable %i (line 6).
-The carry operator 2 is used for a loop variable inserted by the compiler to enforce a memory dependency: since arrays A and B could overlap, we have an additional dataflow path from the store operator, passing through the carry operator 2 and the steer operator 5, to the load operator, in order to enforce that the load in the next iteration waits until the previous store has finished.
-Bisimulation Check
-The first step in WaveCert’s translation validation procedure is to check that the LLVM function is bisimilar to the dataflow program on a canonical schedule of operators, where the operators are executed in the same order as their LLVM instruction counterparts.
-To check a bisimulation, we need to construct a relation between the states of the LLVM and dataflow programs.
+The carry operator 1 is used for the loop induction variable %i (line 5).
+The carry operator 2 is used for a loop variable inserted by the compiler to enforce a memory dependency: since arrays A and B could overlap, we have an additional dataflow path along operators $9\to 2\to 5\to 7$ in order to enforce that the load in the next iteration waits until the previous store has finished.
+Figure 3.
+This plot shows a trace of execution from dataflow cut point 2 (upper left corner).
+The operators fired in this trace are 1, 2, 4, 3, 5, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2).
+At the end of the trace, the configuration matches cut point 2 again.
+In each configuration, we mark the fired operators with orange lightning symbols.
+The emptiness of each channel after firing the operator(s) is represented by dashed vs solid lines.
+Before the firing of an operator, the input channels of the operator should be non-empty and marked with a solid line.
+After firing, the input channels are emptied, and the modified output channels will become non-empty and marked with a solid, orange line.
+The symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked.
+At the top of each configuration, we also indicate the permission constraints added at each step.
+[…]
+Cut point 2 trace.
+Index LLVM Program Point Dataflow Configuration Correspondence
+1 %entry Initial config $\mathsf{Mem}_{\mathrm{llvm}}=\mathsf{Mem}_{\mathrm{df}}$
+[…]
+$\bot$ %end Final config $\mathsf{Mem}_{\mathrm{llvm}}=\mathsf{Mem}_{\mathrm{df}}$
+[…]
+[…]
+Cut points for the example.
+Simulation Check
+The first step in WaveCert’s translation validation procedure is to check that the LLVM and dataflow program are equivalent on a canonical schedule of dataflow operators, where the dataflow operators are executed in the same order as their LLVM instruction counterparts.
+This is done using a simulation relation, or relational invariant, between the states of LLVM and dataflow programs.
 WaveCert constructs this relation by placing corresponding cut points on both the LLVM and the dataflow sides.
-A cut point symbolically describes a set of pairs of LLVM and dataflow configurations satisfying a particular correspondence condition, and a list of cut points together describes the bisimulation relation.
+A cut point symbolically describes a set of pairs of LLVM and dataflow configurations satisfying a correspondence condition; a list of cut points together describes the simulation relation.
 Figure 4 shows the list of cut points for the example.
@@
 For each pair of LLVM and dataflow cut points, WaveCert also infers a list of equations expressing the correspondence of symbolic variables in these cut points.
-To automatically check that the proposed relation in Figure 4 is indeed a bisimulation, we perform symbolic execution from all pairs of cut points (except for $\bot$) until they reach another pair of cut points.
-Symbolic execution (symbolic-execution, ) is similar to concrete execution of programs, except with concrete values replaced by symbolic expressions.
-In this way, we can execute all possible concrete instances of a cut point altogether.
+To automatically check that the proposed relation in Figure 4 is indeed a simulation, we perform symbolic execution (Baldoni et al., 2018) from all pairs of cut points (except for $\bot$) until they reach another pair of cut points.
 On the LLVM side, we first symbolically execute cut point 1, which turns into two symbolic branches: reaching %header after one loop iteration (cut point 2), or failing the loop condition and reaching %end (cut point $\bot$).
@@
 We use this mapping to infer the order in which we fire the dataflow operators, i.e., the canonical schedule.
-For example, from LLVM cut point 2, there are two branches going to cut point 2 and cut point $\bot$.
+For example, from LLVM cut point 2, there are two branches reaching cut point 2 and cut point $\bot$ respectively.
 These two branches have two traces of LLVM instructions (excluding br):
-  • For the first branch to cut point 2: lines 6, 7, 11 - 16;
-  • For the second branch to cut point $\bot$: lines 6, 7.
-For the first branch, the list of LLVM instructions map to dataflow operators 1, 2, 4, 7, 8, 9, 6 (excluding ones that do not have a corresponding operator; operator 2 corresponds to a PHI instruction implicitly inserted during compilation for memory ordering).
+  • For the first branch to cut point 2: lines 5, 6, 9 - 14;
+  • For the second branch to cut point $\bot$: lines 5, 6.
+For the first branch, the LLVM instructions map to dataflow operators 1, 2, 4, 7, 8, 9, 6 (excluding ones that do not have a corresponding operator; operator 2 corresponds to a phi instruction implicitly inserted during compilation for memory ordering).
 LLVM instructions in the second branch map to dataflow operators 1, 2, and 4.
 Therefore, for dataflow cut point 2, we execute these two traces of dataflow operators (and also any steer operator that can be executed), which gives us two dataflow configurations.
-Figure 3 shows the first branch from cut point 2 to itself.
-At the end of the execution, we can see that the configuration matches the cut point 2 configuration we started with, except with potentially different memory configurations and symbolic variables (e.g.
+Figure 3 shows the trace of execution of the first branch from cut point 2 to itself.
+At the end of the execution, we can see that the configuration matches the cut point 2 we started with, except with a different memory state and symbolic variables (e.g.
 i + 1 vs.
@@
 […]
-$\mathrm{mem}_{\mathrm{llvm}}[\texttt{\%B}[\texttt{\%i}]\mapsto\texttt{\%A}[\texttt{\%i}]+1]=\mathrm{mem}_{\mathrm{df}}[\texttt{B}[\texttt{i}]\mapsto\texttt{A}[\texttt{i}]+1]$
-%i + 1=i + 1%i + $=\texttt{i + 1}$ i + 1
-If this check succeeds for all dataflow and LLVM cut points, we soundly conclude that we have obtained a bisimulation between the LLVM program and the dataflow program (under the deterministic canonical schedule).
+[…]
+If this check succeeds for all cut points, we soundly conclude that we have obtained a simulation relation between the LLVM program and the dataflow program on the canonical schedule.
 Confluence Check
-Once we have established a bisimulation between the LLVM function and the canonical schedule of the dataflow program, we check that the dataflow program is confluent: all possible schedules converge to the canonical schedule in the final state (or synchronize with it infinitely often if the program is non-terminating).
+Once we have established a simulation relation between the LLVM program and a canonical schedule of the dataflow program, we check that the dataflow program is confluent: all possible schedules converge to the canonical schedule in the final state (or synchronize with it infinitely often if the program is non-terminating).
+This confluence result ensures that our simulation result for the canonical schedule is fully general for all other schedules.
 […]
-We attach a permission variable (representing a permission token that we do not know a priori) to each value in the channels of dataflow configurations in each symbolic branch.
+In general, as defined in Section 5, we allow write to split into $k$ tokens of read, for a predetermined value of $k$, and we have one permission token for every memory region, such as the array A or B.
+The intuition is that to write to a memory location, an operator needs to have exclusive write ownership of that location and no other operator should have write or read permissions; while to read a memory location, we allow potentially $k$ parallel reads to happen independently.
+If all reads are done, and a write needs to occur, we merge all read tokens into one write token.
+These permission tokens are passed between operators via channels, and operators are not allowed to duplicate or create new tokens (i.e., permission tokens are used linearly).
+We attach permissions to values flowing through the dataflow program, instead of creating separate channels for communicating permission tokens.
+To determine the exact assignment and flow of permission tokens, we first attach a permission variable (representing a permission token that we do not know a priori) to each value in the channels of dataflow configurations in each symbolic branch.
 Figure 3 shows an example of this attachment of permission variables and permission constraints.
@@
 After the first step in which operators 1 and 2 fire, these permission variables are consumed by the two carry operators and we generate fresh permission variables $p_{5},p_{6},p_{7}$ for the new output values.
-We require that the permission variables are used in a linear fashion: operators cannot duplicate or generate permission tokens, instead, they have to be obtained from the inputs.
+We require that the permission variables are used in a linear fashion: operators cannot duplicate or generate permission tokens; instead, they have to be obtained from the inputs.
 So for the first step in Figure 3, we add two constraints $p_{5}+p_{6}\leq p_{1}+p_{3}$ and $p_{7}\leq p_{2}+p_{4}$ to say that the output permissions have be contained in the input permissions.
-Furthermore, if the operator is a load or a store, such as in Figure 3 steps 4 and 6, we require that the suitable permission (read for load, and write for store) should be present in the input permissions (e.g.
+Furthermore, if the operator is a load or a store, such as in Figure 3 steps 4 and 6, we require that the suitable permission (read for load, and write for store) is present in the input permissions (e.g.
 $\texttt{write}\leq p_{13}+p_{17}$ at step 6 in Figure 3).
@@
 If the configuration is matched to a cut point, we add additional constraints to make sure that the assignment of tokens at the cut point is consistent.
-For example in Figure 3, for the last configuration, we would add these additional constraints: $p_{1}=p_{19}$, $p_{2}=p_{18}$, $p_{3}=p_{8}$, $p_{4}=p_{9}$.
-We then solve for the satisfiability of these constraints with respect to the permission algebra ($(\{0,\texttt{read},\texttt{write}\},+,\leq)$ we have defined above.
+For example in Figure 3, since the last configuration is matched again to the same cut point 2, we would add these additional constraints: $p_{1}=p_{19}$, $p_{2}=p_{18}$, $p_{3}=p_{8}$, $p_{4}=p_{9}$.
+We then solve for the satisfiability of these constraints with respect to the permission algebra ($(\{0,\texttt{read}_{1},\texttt{read}_{2},\texttt{write}\},+,\leq)$ that we have defined above.
 If there is a satisfying assignment of the permission variables with permission tokens, then intuitively, the operators in the dataflow program are consistent in memory accesses, and there will not be any data races.
@@
 For the example in Figure 3, a satisfying assignment is $p_{2}=p_{7}=p_{15}=p_{16}=p_{17}=p_{18}=\texttt{write}$ with all other variables set to 00.
-This assignment essentially passes the write permission across the loop of operators 2, 5, 7, 8, 9.
-4. Bisimulation on the Canonical Schedule
-Now, we detail the first half of our translation validation technique, which proves via cut-bisimulation that the LLVM program is equivalent to the dataflow program when restricted to a canonical schedule.
-Given a dataflow program and LLVM program, their semantics are defined by transition systems $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$, respectively, where $\mathcal{I}_{\mathrm{df}}$ and $\mathcal{I}_{\mathrm{llvm}}$ are the sets of initial configurations and $\mathcal{C}_{\mathrm{df}}$ and $\mathcal{C}_{\mathrm{llvm}}$ are the sets of all configurations.
-We use $\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})$ to denote the state of the memory in a dataflow configuration $c_{\mathrm{df}}$, and similarly $\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})$ for the memory in an LLVM configuration $c_{\mathrm{llvm}}$.
+This assignment essentially passes the write permission across the loop of operators $2\to 5\to 7\to 8\to 9\to 2$.
+4. Equivalence on a Canonical Schedule via Cut-Simulation
+We detail the first half of our translation validation technique, which proves via cut-simulation that the LLVM program is equivalent to the dataflow program when restricted to a canonical schedule of dataflow operators.
+Most of our formulation is independent of the source language.
+Hence, instead of using LLVM directly, we assume some sequential and imperative input language called Seq such that 1) Seq is deterministic; 2) Seq has a sequentially consistent memory model (as do dataflow programs); 3) Seq programs operate on fixed-width integer values (as do dataflow programs).
+Throughout this section, we fix a dataflow program and a Seq program, and we denote their operational semantics as transition systems $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{seq}},\mathcal{C}_{\mathrm{seq}},\to_{\mathrm{seq}})$, respectively, where $\mathcal{I}_{\mathrm{df}}$ and $\mathcal{I}_{\mathrm{seq}}$ are the sets of initial configurations, and $\mathcal{C}_{\mathrm{df}}$ and $\mathcal{C}_{\mathrm{seq}}$ are the sets of all configurations.
+We use $\mathsf{Mem}_{\mathrm{df}}(c_{\mathrm{df}})$ to denote the state of the memory in a dataflow configuration $c_{\mathrm{df}}$, and similarly $\mathsf{Mem}_{\mathrm{seq}}(c_{\mathrm{seq}})$ for the memory in a Seq configuration $c_{\mathrm{seq}}$.
 We will omit the subscripts when it is clear from the context.
-For the dataflow semantics, we use $\to^{o}_{\mathrm{df}}$ to denote the (deterministic) firing of a particular operator $o$.
-For a transition relation $\to$, we use $\to^{*}$ for the reflexive-transitive closure of $\to$, and $\to^{+}$ for the transitive closure of $\to$.
+We assume that the memory map is modeled in the same way in both semantics, so that we can directly compare them by $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$.
+Additionally, both the dataflow and Seq programs have input parameters (e.g., the pointer %A in Figure 2(b)); thus, we assume a relation $\sim_{\mathcal{I}}\ \subseteq\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{seq}}$ on initial configurations that guarantees equal values on all input parameters.
+For a transition relation $\to\ \in\{\to_{\mathrm{df}},\to_{\mathrm{seq}}\}$, we use $\to^{*}$ for the reflexive-transitive closure of $\to$, and $\to^{+}$ for the transitive closure of $\to$.
 A configuration $c$ is terminating/final if there exists no configuration $c^{\prime}$ such that $c\to c^{\prime}$.
-One important tool that we use is symbolic execution, which is similar to concrete execution but performed with symbolic expressions in the configuration.
-In the following, we establish some notation for symbolic execution.
+A trace is a finite or infinite sequence of configurations $\tau=c_{1}\dots c_{i}\dots$ such that $c_{i}\to c_{i+1}$ for any $i$.
+A trace $\tau$ is complete if $\tau$ is infinite or the last configuration of $\tau$ is final.
+We define $\mathsf{Traces}_{\mathrm{df}}$ and $\mathsf{Traces}_{\mathrm{seq}}$ to be the set of all traces in the dataflow and Seq semantics, respectively.
+For the dataflow semantics, we use $\to^{o}_{\mathrm{df}}$ to denote the deterministic firing of a particular operator $o$.
+One important tool that we use is symbolic execution, and in the following, we establish some notation for symbolic execution.
 Let $\mathbf{x}$ be a list of variables $\mathbf{x}=(x_{1},\dots,x_{n})$.
-Let $V$ be the set of all bit-vectors of any lengths.
-We use $V(\mathbf{x})$ to denote the set of symbolic expressions constructed from constants in $V$, variables in $\mathbf{x}$, and any bit-vector operation we could perform in our semantics (e.g.
-32-bit zero bit-vector $\texttt{0}_{32}$, or terms like $x_{1}+\texttt{1}_{32}$ are in $V(\mathbf{x})$).
-For $s\in\{\mathrm{df},\mathrm{llvm}\}$, we use $\mathcal{C}_{s}(\mathbf{x})$ to represent the set of symbolic configurations with free variables $\mathbf{x}$ in the semantics $s$.
-A symbolic configuration $c(\mathbf{x})\in\mathcal{C}_{s}(\mathbf{x})$ is similar to a configuration but over symbolic expressions $V(\mathbf{x})$ instead of only $V$ (for example, a symbolic dataflow configuration can have symbolic expressions such as $x_{1}+\texttt{1}_{32}$ in the channels; a symbolic LLVM configuration can have symbolic expressions stored in variables as opposed to only concrete values in $V$).
-If we have a substitution $\sigma:\{x_{1},\dots,x_{n}\}\to V(\mathbf{y})$ from $\mathbf{x}$ to symbolic expressions over some other list of variables $\mathbf{y}$, we denote $c(\sigma)\in\mathcal{C}_{s}(\mathbf{y})$ to be the result of the substitution.
-We also lift the transition relation $\to_{s}$ (for $s\in\{\mathrm{df},\mathrm{llvm}\}$) to symbolic configurations: we use $c(\mathbf{x})\to_{s}c^{\prime}(\mathbf{x})\mid\varphi(\mathbf{x})$ to denote the symbolic execution from $c(\mathbf{x})$ to $c^{\prime}(\mathbf{x})$ with path condition $\varphi(\mathbf{x})$.
-The dataflow transition relation $\to_{\mathrm{df}}$ could be nondeterministic (a configuration can have multiple outgoing transitions), while $\to_{\mathrm{llvm}}$ is deterministic.
-However, for this section, we only care about a canonical schedule for the execution of operators.
-We denote the transition relation restricted to this canonical schedule by $\to_{\mathrm{df\mid canon}}\subseteq\to_{\mathrm{df}}$, and we focus on this deterministic transition for this section.
-The exact meaning of this canonical schedule will be explained in Section 4.2, but essentially, it executes dataflow operators in the same order their corresponding LLVM instructions would execute.
-The equivalence property that we want to check is that for any pair of concrete initial states $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}$ with $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$,
-  • $c_{\mathrm{df}}\to^{*}_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$ for a final state $c^{\prime}_{\mathrm{df}}$ iff $c_{\mathrm{llvm}}\to^{*}_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$ for a final state $c^{\prime}_{\mathrm{llvm}}$, in which case $\mathrm{mem}(c^{\prime}_{\mathrm{df}})=\mathrm{mem}(c^{\prime}_{\mathrm{llvm}})$.
-  • $c_{\mathrm{df}}$ is non-terminating iff $c_{\mathrm{llvm}}$ is non-terminating, and there are infinitely many pairs of states in their traces with equal memory.
-Since the dataflow and LLVM traces could both be unbounded, naive symbolic execution from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{llvm}})$ may not terminate.
-So we need to prove a coinductive invariant about the pairs of traces starting from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{llvm}})$, and such an invariant is a cut-bisimulation relation.
-Definition 0 (Cut-bisimulation).
-Let $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}$ be a binary relation.
-We say that $\mathcal{R}$ is a cut-bisimulation iff for any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$, if $c_{\mathrm{df}}\to_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$ for some $c^{\prime}_{\mathrm{df}}$, then there are $(c^{\prime\prime}_{\mathrm{df}},c^{\prime}_{\mathrm{llvm}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{df}}\to^{*}_{\mathrm{df\mid canon}}c^{\prime\prime}_{\mathrm{df}}$ and $c_{\mathrm{llvm}}\to^{+}_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$.
-Conversely, if $c_{\mathrm{llvm}}\to_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$ for some $c^{\prime}_{\mathrm{llvm}}$, then there are $(c^{\prime}_{\mathrm{df}},c^{\prime\prime}_{\mathrm{llvm}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{llvm}}\to^{*}_{\mathrm{llvm}}c^{\prime\prime}_{\mathrm{llvm}}$ and $c_{\mathrm{df}}\to^{+}_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$.
-This is a weaker variant of the usual notion of bisimulation where we only allow one step before synchronizing instead of allowing one or more steps.
-Work by Kasampalis et al.
-(keq, ) provides more detail about this variant of bisimulation.
-To check equivalence, we look for cut-bisimulations that are memory-synchronizing:
-Definition 0 ().
-A binary relation $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}$ is memory-synchronizing iff
-  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\subset\mathcal{R}$, $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$.
-  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}$ with $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$, $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$.
-  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})$ with $c_{\mathrm{df}},c_{\mathrm{llvm}}$ final and $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$, $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$.
-If there exists a cut-bisimulation $\mathcal{R}$ that is memory synchronizing, then we can conclude via a coinductive argument that the two programs are equivalent.
-In the following subsections, we first describe how we represent a cut-bisimulation relation via cut points (Section 4.1), and how to automatically check that a proposed relation is actually a cut-bisimulation via symbolic execution (Section 4.2).
-Then in Section 4.3, we describe how the cut points are inferred using hints from the compiler.
-4.1. Describing a Cut-Bisimulation via Cut Points
-To finitely describe a proposed cut-bisimulation, we use a list of pairs of dataflow and LLVM cut points.
-A cut point is a pair $P=(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{llvm}}(\mathbf{y}))\mid\varphi(\mathbf{x},\mathbf{y})$ of symbolic dataflow and LLVM configurations constrained by the correspondence condition $\varphi(\mathbf{x},\mathbf{y})$.
-In our case, the condition $\varphi(\mathbf{x},\mathbf{y})$ is a conjunction of equations between variables in $\mathbf{x}$ and $\mathbf{y}$.
-In particular, we require $\mathrm{mem}(c_{\mathrm{df}}(\mathbf{x}))=\mathrm{mem}(c_{\mathrm{llvm}}(\mathbf{y}))$ to be in $\varphi(\mathbf{x},\mathbf{y})$ for any cut point.
-Semantically, each cut point describes the relation $\mathcal{R}(P):=\{(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{llvm}}(\mathbf{y}))\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}\mid\mathbf{x},\mathbf{y}\in V,\varphi(\mathbf{x},\mathbf{y})\}$.
+Let $V$ be the set of all bit-vectors of a predetermined length (for modeling fixed-width integers).
+We use $\mathsf{Terms}(\mathbf{x})$ to denote the set of symbolic expressions constructed from constants in $V$, variables in $\mathbf{x}$, and any bit-vector operation we could perform in our semantics (e.g., a 32-bit zero bit-vector $\texttt{0}_{32}$, or a term like $x_{1}+\texttt{1}_{32}$ are in $\mathsf{Terms}(\mathbf{x})$).
+For $s\in\{\mathrm{df},\mathrm{seq}\}$, we use $\mathcal{C}_{s}(\mathbf{x})$ to represent the set of symbolic configurations with free variables $\mathbf{x}$ in the semantics $s$.
+A symbolic configuration $c(\mathbf{x})\in\mathcal{C}_{s}(\mathbf{x})$ is similar to a configuration but over symbolic expressions $\mathsf{Terms}(\mathbf{x})$ instead of only $V$ (for example, a symbolic dataflow configuration can have symbolic expressions such as $x_{1}+\texttt{1}_{32}$ in the channels).
+If we have a substitution $\sigma:\{x_{1},\dots,x_{n}\}\to\mathsf{Terms}(\mathbf{y})$ from $\mathbf{x}$ to symbolic expressions over some other list of variables $\mathbf{y}$, we denote $c(\sigma)\in\mathcal{C}_{s}(\mathbf{y})$ to be the result of the substitution.
+We use $c(\mathbf{x})\to_{s}c^{\prime}(\mathbf{x})\mid\varphi(\mathbf{x})$ to denote symbolic execution from $c(\mathbf{x})$ to $c^{\prime}(\mathbf{x})$ with path condition $\varphi(\mathbf{x})$.
+A dataflow program does not have a notion of a return value, and the only observable state after finishing execution is the modified global memory.
+Thus we first define when two traces are equivalent based on their memory state.
+Definition 0 (Memory-synchronizing traces).
+Let $\tau_{\mathrm{df}}=c^{1}_{\mathrm{df}}\dots c^{i}_{\mathrm{df}}\dots$ and $\tau_{\mathrm{seq}}=c^{1}_{\mathrm{seq}}\dots c^{j}_{\mathrm{seq}}\dots$ be two complete traces starting from $c^{1}_{\mathrm{df}}$ and $c^{1}_{\mathrm{seq}}$, respectively.
+The traces $\tau_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing iff $\mathsf{Mem}(c^{1}_{\mathrm{df}})=\mathsf{Mem}(c^{1}_{\mathrm{seq}})$, and
+  • If $\tau_{\mathrm{df}}$ is finite, then $\tau_{\mathrm{seq}}$ is finite and $\mathsf{Mem}(c^{\prime}_{\mathrm{df}})=\mathsf{Mem}(c^{\prime}_{\mathrm{seq}})$, where $c^{\prime}_{\mathrm{df}}$ and $c^{\prime}_{\mathrm{seq}}$ are the final states of $\tau_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$, respectively.
+  • If $\tau_{\mathrm{df}}$ is infinite, then $\tau_{\mathrm{seq}}$ is infinite, and for any $n\in\mathbb{N}$, there are $i,j\geq n$ such that $\mathsf{Mem}(c^{i}_{\mathrm{df}})=\mathsf{Mem}(c^{j}_{\mathrm{seq}})$.
+In other words, two complete dataflow and Seq traces are memory-synchronizing if they are both finite and have the same memory state in the initial and final configurations (i.e., both programs terminate and have the same final memory state); or they are both infinite and they synchronize in the same memory state infinitely often.
+Now, we may define program equivalence, which describes when two matching initial configurations for dataflow and Seq exhibit the same observable behaviors:
+Definition 0 (Program equivalence).
+Two programs $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{seq}},\mathcal{C}_{\mathrm{seq}},\to_{\mathrm{seq}})$ are equivalent if for any pair of initial states $c_{\mathrm{df}}\sim_{\mathcal{I}}c_{\mathrm{seq}}$ with $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, if we take any complete dataflow trace $\tau_{\mathrm{df}}=c_{\mathrm{df}}\dots$ and the unique complete Seq trace $\tau_{\mathrm{seq}}=c_{\mathrm{seq}}\dots$ (unique since Seq is deterministic), then $\tau_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing.
+Intuitively, the dataflow program and Seq program are equivalent if any pair of complete traces from corresponding initial states are memory-synchronizing; i.e., they either both terminate with the same memory state, or both run forever and synchronize infinitely often.
+Our final goal is to achieve Definition 2.
+We will do so by first verifying a weaker property of equivalence along a particular, deterministic canonical schedule of the dataflow program.
+Then, in Section 5, we show that this canonical schedule suffices to prove equivalence for all possible dataflow schedules.
+First, we introduce canonical schedules, which determinize dataflow programs using the information contained in Seq traces:
+Definition 0 (Canonical schedule).
+A canonical schedule is a function $\mathsf{Canon}:\mathcal{C}_{\mathrm{df}}\times\mathsf{Traces}_{\mathrm{seq}}\to\mathsf{Traces}_{\mathrm{df}}$ that selects a dataflow trace using a Seq trace, such that
+  • For any $c_{\mathrm{df}}\in\mathcal{C}_{\mathrm{df}}$ and any non-empty Seq trace $\tau_{\mathrm{seq}}$, $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is a dataflow trace starting in $c_{\mathrm{df}}$.
+  • If $\tau_{\mathrm{seq}}$ is finite, then $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is finite.
+  • For any $c_{\mathrm{df}}\in\mathcal{C}_{\mathrm{df}}$, if $\tau^{\prime}_{\mathrm{seq}}$ is a prefix of $\tau_{\mathrm{seq}}$, then $\mathsf{Canon}(c_{\mathrm{df}},\tau^{\prime}_{\mathrm{seq}})$ is a prefix of $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$.
+We then specialize program equivalence (Definition 2) by restricting our attention to those canonical dataflow schedules:
+Definition 0 (Program equivalence on a canonical schedule).
+Let $\mathsf{Canon}$ be a canonical schedule.
+Two programs $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{seq}},\mathcal{C}_{\mathrm{seq}},\to_{\mathrm{seq}})$ are equivalent on the canonical schedule $\mathsf{Canon}$ if for any pair of initial states $c_{\mathrm{df}}\sim_{\mathcal{I}}c_{\mathrm{seq}}$ with $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, if we take the unique complete Seq trace $\tau_{\mathrm{seq}}=c_{\mathrm{seq}}\dots$, then $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is a complete dataflow trace, and $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing.
+That is, instead of enforcing that all complete dataflow traces are memory-synchronizing to the complete Seq trace, we only require that there exists one such trace selected by some canonical schedule $\mathsf{Canon}$.
+Later in Section 5, we separately verify the confluence property (Definition 1), so that Definition 1 and Definition 4 together imply Definition 2 (Proposition 2).
+Canonical Schedules for RipTide
+[…]
+$\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{llvm}}):=\begin{cases}\epsilon&\text{If }\tau_{\mathrm{llvm}}=\epsilon\\ \tau_{\mathrm{df}}\cdot\mathsf{Canon}(c^{\prime}_{\mathrm{df}},\tau^{\prime}_{\mathrm{llvm}})&\text{If }\tau_{\mathrm{llvm}}=c_{\mathrm{llvm}}\cdot\tau^{\prime}_{\mathrm{llvm}}\end{cases}$
+[…]
+In the following sections, we outline the steps to verify that the dataflow and Seq programs satisfy Definition 4.
+In Section 4.1, we first describe our general method of using a coinductive invariant called cut-simulation that implies the property in Definition 4.
+In Section 4.2, we show how cut-simulation can be symbolically represented by cut points.
+In Section 4.3, we present the algorithm to check that a set of cut points does form a cut-simulation.
+Finally in Section 4.4, we describe our heuristics to infer cut points for the compilation from LLVM to dataflow programs.
+4.1. Proving Equivalence on the Canonical Schedule via Cut-Simulation
+Traces from $\mathcal{I}_{\mathrm{df}}$ and $\mathcal{I}_{\mathrm{seq}}$ are unbounded in general, so to prove the equivalence on a canonical schedule (Definition 4), we need to verify a coinductive invariant about the pairs of traces starting from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{seq}})$, and such an invariant is a cut-simulation relation (Kasampalis et al., 2021).
+Definition 0 (Cut-simulation).
+Let $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{seq}}$ be a binary relation.
+We say that $\mathcal{R}$ is a cut-simulation between $(\mathcal{I}_{\mathrm{seq}},\mathcal{C}_{\mathrm{seq}},\to_{\mathrm{seq}})$ and $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ iff for any $(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{R}$, if $c_{\mathrm{seq}}\to_{\mathrm{seq}}c^{\prime}_{\mathrm{seq}}$ for some $c^{\prime}_{\mathrm{seq}}$, then there are $(c^{\prime}_{\mathrm{df}},c^{\prime\prime}_{\mathrm{seq}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{seq}}\to^{*}_{\mathrm{seq}}c^{\prime\prime}_{\mathrm{seq}}$ and $c_{\mathrm{df}}\to^{+}_{\mathrm{df}}c^{\prime}_{\mathrm{df}}$.
+Intuitively, a cut-simulation states that if the two transition systems are synchronized in $\mathcal{R}$, then if we can take a step on the Seq side, we can step the two transition systems further to re-synchronize them in $\mathcal{R}$.
+Note that in our setting, we are verifying that the dataflow program simulates the Seq program (i.e., the target program simulates the source program), instead of the other way around which is common in compiler verification work.
+This is because our target dataflow program has more nondeterministic behavior than the source Seq program, whereas in many other cases, the source program has more behavior (such as undefined behavior in C) than the concrete target program.
+To satisfy our desired equivalence properties, we look for cut-simulations that are memory-synchronizing:
+Definition 0.
+A binary relation $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{seq}}$ is memory-synchronizing iff
+  • For any $(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{R}$, $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$.
+  • For any initial configurations $c_{\mathrm{df}}\sim_{\mathcal{I}}c_{\mathrm{seq}}$ with $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, $(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{R}$.
+  • For any $(c_{\mathrm{df}},c_{\mathrm{seq}})$ with $c_{\mathrm{df}},c_{\mathrm{seq}}$ final and $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, $(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{R}$.
+The second and third conditions enforce that $\mathcal{R}$ should include all pairs of corresponding initial states, as well as final states; thus $\mathcal{R}$ covers all complete traces.
+If there exists a memory-synchronizing cut-simulation $\mathcal{R}$, then we can conclude via a coinductive argument that the two programs are equivalent on a canonical schedule (Definition 4).
+4.2. Describing a Cut-Simulation via Cut Points
+To finitely describe a proposed cut-simulation, we use a list of pairs of dataflow and Seq cut points.
+A cut point is a pair $P=(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{seq}}(\mathbf{y}))\mid\varphi(\mathbf{x},\mathbf{y})$ of symbolic dataflow and Seq configurations constrained by the correspondence condition $\varphi(\mathbf{x},\mathbf{y})$.
+In our case, the condition $\varphi(\mathbf{x},\mathbf{y})$ is a conjunction of equalities between variables in $\mathbf{x}$ and $\mathbf{y}$.
+Semantically, each cut point describes the binary relation $\mathcal{R}(P):=\{(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{seq}}(\mathbf{y}))\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{seq}}\mid\mathbf{x},\mathbf{y}\in V,\varphi(\mathbf{x},\mathbf{y})\}$.
 For a list of $n$ cut points
-$\{P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\}_{i=1}^{n},$
+$\{P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{seq}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\}_{i=1}^{n},$
 the relation they describe is the union of the subrelation each describes: $\mathcal{R}(P^{1},\dots,P^{n}):=\bigcup^{n}_{i=1}\mathcal{R}(P^{i})$.
-To ensure that the cut-bisimulation we propose satisfies the property in Definition 2, we also require that the first cut point must be exactly the pairs of initial configurations with equal memory.
-That is $\mathcal{R}(P^{1})=\{(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}\mid\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})=\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})\}$.
-Furthermore, we require that the last cut point $P^{n}$ must be the pairs of final configurations with equal memory, that is $\mathcal{R}(P^{n})=\{(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}\mid c_{\mathrm{df}},c_{\mathrm{llvm}}\text{ final},\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})=\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})\}$.
+To ensure that the cut-simulation we propose satisfies the three conditions in Definition 6, we only construct cut points such that:
+  • For any cut point $P^{i}$, the correspondence constraint $\varphi^{i}(\mathbf{x}^{i},\mathbf{y}^{i})$ implies equal memory $\mathsf{Mem}(c^{i}_{\mathrm{df}}(\mathbf{x}^{i}))=\mathsf{Mem}(c^{i}_{\mathrm{seq}}(\mathbf{y}^{i}))$.
+  • The first cut point exactly represents the pairs of equivalent, initial configurations with equal memory; i.e., $\mathcal{R}(P^{1})=\{(c_{\mathrm{df}},c_{\mathrm{seq}})\in\ \sim_{\mathcal{I}}\ \mid\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})\}$.
+  • The last cut point $P^{n}$ exactly represents the pairs of final configurations with equal memory; i.e., $\mathcal{R}(P^{n})=\{(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{seq}}\mid c_{\mathrm{df}},c_{\mathrm{seq}}\text{ final},\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})\}$.
 Then $\mathcal{R}(P^{1},\dots,P^{n})$ satisfies Definition 6.
-4.2. Checking Cut-Bisimulation
-Now that we have a way to symbolically describe a relation $\mathcal{R}(P^{1},\dots,P^{n})$ with cut points satisfying Definition 2, we need to check that it is indeed a cut-bisimulation.
+4.3. Checking Cut-Simulation
+Now that we have a way to symbolically describe a relation $\mathcal{R}(P^{1},\dots,P^{n})$ with cut points satisfying Definition 6, we need to check that it is indeed a cut-simulation.
 To achieve this, we perform symbolic execution from each cut point and try to reach another cut point.
-Suppose we are checking cut point $P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$.
-We perform symbolic execution from $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ and $c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})$ until all of their symbolic branches hit another cut point.
-For the LLVM side, we get a list of $k_{i}$ symbolic branches with path conditions denoted by $\psi$:
-$c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})\to^{+}_{\mathrm{llvm}}c^{i,1}_{\mathrm{llvm}}(\mathbf{y^{i}})\mid\psi_{\mathrm{llvm},1}(\mathbf{y^{i}})$
+Suppose we are checking cut point $P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{seq}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$.
+We perform the following three steps.
+Step 1. Symbolic execution of Seq cut points
+We perform symbolic execution from $c^{i}_{\mathrm{seq}}(\mathbf{y^{i}})$ and get a list of $k_{i}$ symbolic branches with path conditions denoted by $\psi$:
+$c^{i}_{\mathrm{seq}}(\mathbf{y^{i}})\to^{+}_{\mathrm{seq}}c^{i,1}_{\mathrm{seq}}(\mathbf{y^{i}})\mid\psi_{\mathrm{seq},1}(\mathbf{y^{i}})$
 $\dots$
-$c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})\to^{+}_{\mathrm{llvm}}c^{i,k_{i}}_{\mathrm{llvm}}(\mathbf{y^{i}})\mid\psi_{\mathrm{llvm},k_{i}}(\mathbf{y^{i}})$
-where each right hand side $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})$ for $l\in\{1,\dots,k_{i}\}$ is an instance $c^{j}_{\mathrm{llvm}}(\sigma)$ of some other LLVM cut point $c^{j}_{\mathrm{llvm}}(\mathbf{y^{j}})$ with substitution $\sigma:\mathbf{y^{j}}\to V(\mathbf{y^{i}})$.
-Furthermore, since our LLVM semantics is deterministic, the path conditions should be disjoint, i.e., for any $l_{1},l_{2}\in\{1,\dots,k_{i}\}$, if $l_{1}\neq l_{2}$, $\psi_{\mathrm{llvm},l_{1}}(\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l_{2}}(\mathbf{y^{i}})$ is unsatisfiable.
-For the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, since the dataflow semantics is nondeterministic, we try to replicate the execution of LLVM branches on the dataflow side.
-For each LLVM branch $l\in\{1,\dots,k_{i}\}$, we can get a trace of LLVM instructions that are executed from $c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})$ to $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})$.
-Using hints from the compiler, we know that a subset of these LLVM instructions map to a list of dataflow operators $(o^{l}_{1},\dots,o^{l}_{n_{l}})$.
-For each LLVM branch, to produce the corresponding dataflow branch, we symbolically execute the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ on the schedule $(o^{l}_{1},\dots,o^{l}_{n_{l}})$ with the path condition $\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$:
-$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{o^{l}_{1}*}_{\mathrm{df}}\dots\to^{o^{l}_{n_{l}}*}_{\mathrm{df}}c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}})\mid\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$
-where each $\to^{o*}_{\mathrm{df}}$ denotes the dataflow transition by first firing the operator $o$, and then firing any available steer operator until none is fireable anymore.
-The canonical schedule $\to_{\mathrm{df\mid canon}}$ is defined by this order of execution of operators.
-Therefore, by replicating the execution of the LLVM branches from cut point $i$, we get a list of dataflow branches:
-(1) $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df\mid canon}}c^{i,1}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},1}(\mathbf{x^{i}})$
+$c^{i}_{\mathrm{seq}}(\mathbf{y^{i}})\to^{+}_{\mathrm{seq}}c^{i,k_{i}}_{\mathrm{seq}}(\mathbf{y^{i}})\mid\psi_{\mathrm{seq},k_{i}}(\mathbf{y^{i}})$
+where each right hand side $c^{i,l}_{\mathrm{seq}}(\mathbf{y^{i}})$ for $l\in\{1,\dots,k_{i}\}$ is an instance $c^{j}_{\mathrm{seq}}(\sigma)$ of some Seq cut point $c^{j}_{\mathrm{seq}}(\mathbf{y^{j}})$ with substitution $\sigma:\mathbf{y^{j}}\to\mathsf{Terms}(\mathbf{y^{i}})$.
+That is, at the end of each branch, the configuration is matched to some cut point.
+Note that since the initial symbolic configuration $c^{i}_{\mathrm{seq}}(\mathbf{y^{i}})$ does not have any constraints, the final path conditions should cover all cases of $\mathbf{y^{i}}$; in other words, $\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{seq},l}(\mathbf{y^{i}})$ is valid.
+Step 2. Symbolic execution of dataflow cut points on the canonical schedule
+For the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, since the dataflow semantics is nondeterministic, the choice of operators to fire is unclear.
+Therefore, we follow a canonical schedule $\mathsf{Canon}$ (Definition 3) to replicate the execution of Seq branches on the dataflow side.
+For each Seq branch $l\in\{1,\dots,k_{i}\}$, we have a Seq trace $\tau_{\mathrm{seq}}=c^{i}_{\mathrm{seq}}(\mathbf{y}^{i})\dots c^{i,l}_{\mathrm{seq}}(\mathbf{y}^{i})$.
+By mapping this trace through the canonical schedule $\tau_{\mathrm{df}}:=\mathsf{Canon}(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),\tau_{\mathrm{seq}})$, we get a list of dataflow operators $(o^{l}_{1},\dots,o^{l}_{n_{l}})$ fired in $\tau_{\mathrm{df}}$.
+Then we symbolically execute the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ on the schedule $(o^{l}_{1},\dots,o^{l}_{n_{l}})$ with the path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}}):=\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{seq},l}(\mathbf{y^{i}})$:
+$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{o^{l}_{1}}_{\mathrm{df}}\dots\to^{o^{l}_{n_{l}}}_{\mathrm{df}}c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},l}(\mathbf{x^{i}})$
+The new path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}})$ amounts to replacing variables in the Seq path condition $\psi_{\mathrm{seq},l}(\mathbf{y^{i}})$ with corresponding variables in the correspondence constraint $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$.
+With the additional path condition imposed, the symbolic execution is not likely to branch.
+If it does, the cut-simulation check will fail; in which case, a more precise heuristic is needed for the canonical schedule.
+By doing this for each Seq branch, we get a list of dataflow branches:
+(1) $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df}}c^{i,1}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},1}(\mathbf{x^{i}})$
 $\dots$
-$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df\mid canon}}c^{i,k_{i}}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},k_{i}}(\mathbf{x^{i}})$
-where each path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}}):=\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$.
-We check that the dataflow symbolic execution from cut point $i$ has covered all concrete cases, i.e., all possible cases of $\mathbf{x^{i}}$, by checking the unsatisfiability of $\neg\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{df},l}(\mathbf{x^{i}})$.
-Finally, we check that each pair of dataflow and LLVM branches $(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}}))$ matches some target cut point $P^{j}=(c^{j}_{\mathrm{df}}(\mathbf{x^{j}}),c^{j}_{\mathrm{llvm}}(\mathbf{y^{j}}))\mid\varphi^{j}(\mathbf{x^{j}},\mathbf{y^{j}})$:
-$\{(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})\}\subseteq\mathcal{R}(P^{j})$
-or equivalently, there are substitutions $\sigma^{i,l}_{\mathrm{df}}:\mathbf{x^{j}}\to V(\mathbf{x^{i}}),\sigma^{i,l}_{\mathrm{llvm}}:\mathbf{y^{j}}\to V(\mathbf{y^{i}})$ such that the following conditions hold:
+$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df}}c^{i,k_{i}}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},k_{i}}(\mathbf{x^{i}})$
+In our case, these branches should cover all possible values of $\mathbf{x}^{i}$ (i.e., $\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{df},l}(\mathbf{x^{i}})$ is valid), since in the path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}})$, the correspondence constraint $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$ only contains a conjunction of equalities between variables in $\mathbf{x^{i}}$ and $\mathbf{y^{i}}$, and the union of all Seq path conditions $\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{seq},l}(\mathbf{y^{i}})$ is valid.
+In general, however, if the correspondence constraint $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$ is more complex (e.g.
+containing formulas such as $x_{1}=2y_{2}$), we may need to additionally verify that $\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{df},l}(\mathbf{x^{i}})$ is valid to ensure that all possible concrete configurations represented by $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ is covered in the symbolic execution.
+Step 3. Cut point subsumption
+Finally, we check that each pair of dataflow and Seq branches $(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{seq}}(\mathbf{y^{i}}))$ is contained in some target cut point $P^{j}=(c^{j}_{\mathrm{df}}(\mathbf{x^{j}}),c^{j}_{\mathrm{seq}}(\mathbf{y^{j}}))\mid\varphi^{j}(\mathbf{x^{j}},\mathbf{y^{j}})$:
+(2) $\{(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{seq}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{seq},l}(\mathbf{y^{i}})\}\subseteq\mathcal{R}(P^{j})$
+or equivalently, there are substitutions $\sigma^{i,l}_{\mathrm{df}}:\mathbf{x^{j}}\to\mathsf{Terms}(\mathbf{x^{i}}),\sigma^{i,l}_{\mathrm{seq}}:\mathbf{y^{j}}\to\mathsf{Terms}(\mathbf{y^{i}})$ such that the following conditions hold:
   • The dataflow branch is an instance of the target dataflow cut point: $c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}})=c^{j}_{\mathrm{df}}(\sigma^{i,l}_{\mathrm{df}})$.
-  • The LLVM branch is an instance of the target LLVM cut point: $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})=c^{j}_{\mathrm{llvm}}(\sigma^{i,l}_{\mathrm{llvm}})$.
+  • The Seq branch is an instance of the target Seq cut point: $c^{i,l}_{\mathrm{seq}}(\mathbf{y^{i}})=c^{j}_{\mathrm{seq}}(\sigma^{i,l}_{\mathrm{seq}})$.
   • The correspondence condition at the target cut point is satisfied given the source cut point correspondence and the path conditions:
-    $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})\rightarrow\varphi^{j}(\sigma^{i,l}_{\mathrm{df}},\sigma^{i,l}_{\mathrm{llvm}})$
-If the check described in this section is passed for all cut points $i$ and all branches $1,\dots,k_{i}$, we can conclude that $\mathcal{R}(P^{1},\dots,P^{n})$ is indeed a cut-bisimulation.
+    $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{seq},l}(\mathbf{y^{i}})\rightarrow\varphi^{j}(\sigma^{i,l}_{\mathrm{df}},\sigma^{i,l}_{\mathrm{seq}})$
+If the steps in this section are passed for all cut points $i$, we can conclude that $\mathcal{R}(P^{1},\dots,P^{n})$ is indeed a cut-simulation.
 4.4. Heuristics to Infer Cut Points
-The final step is to actually infer the LLVM and dataflow cut points.
-In general, program equivalence for Turing-complete models like LLVM and dataflow programs is undecidable (rogers-87, ), so our best hope is to have good heuristics to infer the cut points for the specific input/output programs of the compiler.
+In this section, we describe our heuristics to infer cut points specific to LLVM and dataflow programs.
+In general, program equivalence for Turing-complete models like LLVM and dataflow programs is undecidable, so our best hope is to have good heuristics to infer the cut points for the specific input/output programs of the compiler.
 The placement of cut points is important for the symbolic execution to terminate.
@@
 But the symbolic execution will try to simulate all possible traces together, which are unbounded, and the check may not terminate.
-To ensure termination, on the LLVM side, we place a cut point at each of the following program points:
-  • The entry of the function (i.e. the initial configuration);
-  • The back edge of each loop header block; and
-  • The exit program point.
+To ensure termination, on the LLVM side, we place a cut point at the entry of the function (i.e., the initial configuration), the back edge of each loop header block, and the exit program point.
 For the dataflow program, we need to place the first cut point at the initial configuration of the dataflow program, and the last cut point at the final configuration.
@@
 However, due to the lack of control-flow in the dataflow program, the form of the corresponding dataflow configurations at these “program points” is hard to infer statically.
-Instead of generating the dataflow cut points ahead of time, we dynamically infer them during symbolic execution of dataflow branches.
+Instead of generating the dataflow cut points ahead of time, we dynamically infer them during the symbolic execution of dataflow branches.
 For example, after mirroring the LLVM branch execution in a dataflow branch such as in Equation 1, if the dataflow cut point corresponding to the LLVM cut point has not been inferred yet, we generalize the symbolic configuration on the RHS of the dataflow branch Equation 1 to a dataflow cut point by replacing symbolic expressions with fresh variables.
@@
 Note that we do not need these heuristics to be sound.
-As long as the cut points for the initial and final dataflow/LLVM configurations are generated correctly and the cut-bisimulation check in Section 4.2 succeeds, the soundness of the cut-bisimulation follows.
-While we do not prove these heuristics correct, they work well in practice and cover all test cases in our evaluation (Section 6).
+As long as the cut points for the initial and final dataflow/LLVM configurations are generated correctly and the cut-simulation check in Section 4.3 succeeds, the soundness of the cut-simulation follows.
+While we do not prove that these heuristics are complete, they work well in practice and cover all test cases in our evaluation (Section 6).
+If the heuristics generate incorrect cut points, the simulation check would still terminate but fail.
+To see this, on the LLVM side, our placement of the cut points guarantees termination because we place one at each loop header.
+On the dataflow side, since the execution is mirroring the LLVM execution and no additional search is done, the algorithm would terminate but may fail to match the resulting dataflow configuration against the cut point.
 5. Confluence Checking using Linear Permission Tokens
-After checking cut-bisimilarity between the LLVM program and the dataflow program on the canonical schedule, we still need to prove that this canonical schedule is general.
+After checking the equivalence between the LLVM and dataflow programs on a canonical dataflow schedule, we still need to prove that this canonical schedule is general.
 Indeed, since the dataflow program is asynchronous and nondeterministic, other choices of schedule may lead to data races and inconsistent final states.
-Therefore, in addition to finding a bisimulation on the canonical schedule, we want to prove a confluence property on the dataflow program:
-Definition 0 ().
+Therefore, in addition to checking the equivalence on a canonical schedule, we want to prove a confluence property on the dataflow program:
+Definition 0.
 We say $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is (ground) confluent iff for any initial configuration $c\in\mathcal{I}_{\mathrm{df}}$, if $c\to^{*}_{\mathrm{df}}c_{1}$ and $c\to^{*}_{\mathrm{df}}c_{2}$, then there is $c^{\prime}\in\mathcal{C}_{\mathrm{df}}$ with $c_{1}\to^{*}_{\mathrm{df}}c^{\prime}$ and $c_{2}\to^{*}_{\mathrm{df}}c^{\prime}$.
+If the pair of dataflow and LLVM programs satisfies both Definition 4 and Definition 1, then we can conclude the desired equivalence property of Definition 2.
+Proposition 0.
+If $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$ are equivalent on a canonical schedule (Definition 4) and $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is confluent, then $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$ are equivalent (Definition 2).
+Proof Sketch.
+Assume $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$ are equivalent on some canonical schedule $\mathsf{Canon}$ and $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is confluent.
+Let $c_{\mathrm{df}}\sim_{\mathcal{I}}c_{\mathrm{seq}}$ be any pair of corresponding initial states with $\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})$, and let $\tau_{\mathrm{seq}}=c_{\mathrm{seq}}\dots$ be the unique complete trace.
+Let $\tau^{\prime}_{\mathrm{df}}=c_{\mathrm{df}}\dots$ be any complete trace from $c_{\mathrm{df}}$.
+We need to show that $\tau^{\prime}_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing.
+By equivalence on $\mathsf{Canon}$, we know that $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is a complete trace and $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing.
+We have two cases:
+  • If $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is finite, then by confluence, $\tau^{\prime}_{\mathrm{df}}$ is finite and has the same final state as $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ (since $\tau^{\prime}_{\mathrm{df}}$ is complete).
+  • If $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ is infinite, then by confluence, $\mathsf{Canon}(c_{\mathrm{df}},\tau_{\mathrm{seq}})$ and $\tau^{\prime}_{\mathrm{df}}$ synchronize infinitely often.
+Therefore $\tau^{\prime}_{\mathrm{df}}$ and $\tau_{\mathrm{seq}}$ are memory-synchronizing. ∎
 In this section, we present a technique using what we call linear permission tokens to ensure that the dataflow program is confluent.
-Linear permission tokens are, in essence, a dynamic variant of the fractional permissions studied in the concurrent separation logic (fractional-perm, ; concurrent-separation-logic, ).
-In our case, their linear usage is dynamically enforced and the exact placement of permission tokens is resolved after symbolic execution.
-We associate each value present in a dataflow execution trace with a permission token in the form of $\texttt{write}\ l$, $\texttt{read}\ l$, or any formal sum of them, for any memory location $l$.
-These tokens are used by dataflow operators in a linear fashion: they cannot be replicated, and they are disjoint in the initial state.
-A write token can be split into $n$ many read tokens (for a predetermined parameter $n$), and inversely, $n$ of such read tokens can be combined to get back a write token.
-As shown later in this section, if there is a consistent assignment of these linear permission tokens, we can conclude that there are no data races possible in any schedule, and thus all final states have to be the same.
-In Section 5.1, we prove that the technique works in a bounded setting for a concrete trace $\tau$: if $\tau$ has a consistent assignment of permission tokens, then for any other trace $\tau^{\prime}$ bounded by $\tau$ (meaning that for each operator, $\tau^{\prime}$ executes it for an equal or smaller number of times than $\tau$), $\tau^{\prime}$ will converge to the final configuration of $\tau$.
-Then in Section 5.2, we explain how this bounded confluence check can be extended to a full confluence check for the entire dataflow program by using the cut points that we have placed for the dataflow program.
+Linear permission tokens are, in essence, a dynamic variant of fractional permissions (Boyland, 2013).
+In our case, the exact placement of permission tokens is resolved dynamically via symbolic execution.
+We associate each value flowing through a dataflow program with a permission token in the form of $\texttt{write}\ l$, $\texttt{read}\ l$, or any formal sum of them, for any memory location $l$.
+These tokens are used by dataflow operators in a linear fashion: they cannot be replicated or generated, and they are disjoint in the initial state.
+To support independent parallel reads and enforce that writes must be exclusive, a write token can be split into $k$ many read tokens (for a predetermined parameter $k$), and inversely, $k$ of such read tokens can be merged to get back a write token.
+We show in this section that if there is a consistent assignment of these linear permission tokens to values, then we can conclude that the execution of dataflow program will be confluent.
+This is done in two steps:
+  • In Section 5.1, we prove that the technique works in a bounded setting for a concrete trace $\tau$: if $\tau$ has a consistent assignment of permission tokens, $\tau^{\prime}$ is another trace from the same initial configuration, and $\tau^{\prime}$ is bounded by $\tau$ (meaning that for each operator $o$, the number of times $o$ fired in $\tau^{\prime}$ is less than or equal to the number of times $o$ fired in $\tau$), then there is a valid trace from the final configuration of $\tau^{\prime}$ to the final configuration of $\tau$; i.e., $\tau^{\prime}$ will converge to the final configuration of $\tau$.
+  • In Section 5.2, we explain how this bounded confluence check can be extended to a full confluence check for the entire dataflow program by using the cut points that we have placed for the dataflow program in Section 4.
+Finally in Section 5.3, we show that even though we assume the channels in a dataflow program are unbounded, the cut-simulation and confluence checks allow us to extend the verification results to a bounded-channel model of dataflow programs.
 5.1. Confluence on Bounded Traces
-In this section, we show that if we have a $n$-step trace $c_{1}\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}$ (where $o_{1},\dots,o_{n}$ are the operators fired in each step) with a consistent assignment of permission tokens to each value, then for any other trace $c_{1}\to^{o^{\prime}_{1}}_{\mathrm{df}}\dots\to^{o^{\prime}_{m}}_{\mathrm{df}}c_{m+1}$ (not necessarily with a consistent assignment of permission tokens) such that the multiset $\{o^{\prime}_{1},\dots,o^{\prime}_{m}\}\subseteq\{o_{1},\dots,o_{n}\}$, we have $c_{m+1}\to^{*}_{\mathrm{df}}c_{n+1}$.
+In this section, we show that if we have a finite trace $c_{1}\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}$ (where $o_{1},\dots,o_{n}$ are the operators fired in each step) with a consistent assignment of permission tokens to each value in each configuration, then for any other trace $c_{1}\to^{o^{\prime}_{1}}_{\mathrm{df}}\dots\to^{o^{\prime}_{m}}_{\mathrm{df}}c_{m+1}$ (not necessarily with a consistent assignment of permission tokens) such that the multiset $\{o^{\prime}_{1},\dots,o^{\prime}_{m}\}\subseteq\{o_{1},\dots,o_{n}\}$, we have $c_{m+1}\to^{*}_{\mathrm{df}}c_{n+1}$.
 Let us first describe the permission tokens and their assignment to values. The permission tokens are drawn from the permission algebra defined below.
@@
 Let $L$ be a set of memory locations and $k$ be a positive integer. The permission algebra is a partial algebraic structure $(P(L,k),0,+,\leq)$ where
-  • $(P(L,k),0,+)$ is the term algebra of the formal sums $\texttt{read}\ l_{1}+\dots+\texttt{read}\ l_{n}$, with $l_{1},\dots,l_{n}\in L$, modulo the commutativity and associativity of $+$, and we require that for a location $l$, at most $k$ copies of $\texttt{read}\ l$ is present in the formal sum.
-We call $p+q$ the disjoint union of $p,q\in P(L,k)$.
+  • $P(L,k)$ is the set of permissions such that $(P(L,k),0,+)$ forms a partially commutative monoid.
+It consists of formal sums $\texttt{read}\ l_{1}+\dots+\texttt{read}\ l_{n}$, with $l_{1},\dots,l_{n}\in L$, modulo the commutativity and associativity of $+$.
+We require that for a location $l$, at most $k$ copies of $\texttt{read}\ l$ are present in the formal sum.
+We denote $N(l,p)$ to be the number of occurrences of $\texttt{read}\ l$ in $p$.
+We call $p+q$ the (disjoint) sum of $p,q\in P(L,k)$.
+In particular, for any $l\in L$, $p+q$ is defined iff for all $l$, $N(l,p)+N(l,q)\leq k$.
     We also denote
@@
   • $P(L,k)$ is partially ordered by $\leq$ with $p\leq q$ (reads $q$ contains $p$) iff for any $l$, the number of occurrences of $\texttt{read}\ l$ in $p$ is less than or equal to that in $q$.
-In Boyland’s terminology (fractional-perm, ), $P(L,k)$ is a finite token model of size $k$ of fractional permissions. For example, in $P(\{\texttt{A},\texttt{B}\},2)$, we have that
+Now we define when two or more permissions are considered disjoint, which is crucial for enforcing linearity.
+Definition 0.
+[…]
+For example, in $P(\{\texttt{A},\texttt{B}\},2)$, we have that
 […]
 […]
-Definition 0 ().
-We say that the permissions $p_{1},\dots,p_{n}\in P(L,k)$ are disjoint iff for any $i\neq j$, $p_{i}+p_{j}$ is defined.
-Now let us define an assignment of permission tokens to values in a configuration.
+Intuitively, the reason to allow a write to split into $k$ copies of read is to allow the following two valid scenarios:
+  • A store operator can write to a memory location if it has exclusive write permission to the location (i.e., no other operator has a write or read permission to that memory location);
+  • A load operator can read from a memory location if it has a read permission and all other currently existing permissions in the system are also read.
+Furthermore, once (at most $k$) parallel reads are performed, a store operator can perform a memory write by reclaiming all existing read permissions and merging them into a write; hence we define write as the sum of $k$ copies of read.
+Now let us define the assignment of permission tokens to values in a configuration.
 Let $E$ denote the set of channel names.
@@
 Let $c\in\mathcal{C}_{\mathrm{df}}$ be a dataflow configuration.
-A permission augmentation of $c$ is a partial map $t:E\times\mathbb{N}\to P(L,k)$ such that for any $e\in E$, $t(e,n)$ is defined iff $n\in\{1,\dots,|\mathrm{Channel}(e)|\}$.
+A permission augmentation of $c$ is a partial map $t:E\times\mathbb{N}\rightharpoonup P(L,k)$ such that for any $e\in E$, $t(e,n)$ is defined iff $n\in\{1,\dots,|\mathrm{Channel}(c,e)|\}$.
 We call $(c,t)$ a permission-augmented configuration.
 Intuitively, $t(e,n)$ is the permission we attach to the $n$-th value in the channel $e$.
-In the following, we define when such a permission augmentation is consistent (i.e., the use of permissions is linear).
+Permission augmentations defined above can assign arbitrary permissions to values.
+However, to ensure race-freedom and confluence, we need to restrict them so that, for example, two operators cannot share a write token.
+In the following definitions, we define when a configuration, a transition, and a trace are consistently augmented with permissions, meaning that permissions are used linearly.
 Definition 0 (Consistent Augmentation).
@@
 Let $c\to^{o}_{\mathrm{df}}c^{\prime}$ be a transition.
-Let $t,t^{\prime}$ be two permission augmentations for $c,c^{\prime}$, respectively.
+Let $t,t^{\prime}$ be two consistent permission augmentations for $c,c^{\prime}$, respectively.
 Let $p_{1},\dots,p_{n}$ be the permissions in $t$ attached to input values of $o$ and $q_{1},\dots,q_{m}$ be the permissions in $t^{\prime}$ attached to output values of $o$.
 We say $(c,t)\to^{o}_{\mathrm{df}}(c^{\prime},t^{\prime})$ is a consistent (permission-augmented) transition iff:
-  • The $t$ and $t^{\prime}$ have the same assignment of permissions except for the channels changed by the operator $o$.
+  • $(c,t)$ and $(c^{\prime},t^{\prime})$ are consistent.
+  • $t(e,n)=t^{\prime}(e,n)$ for all channel $e$ and position $n$ except for those changed by $o$.
   • (Linearity) $q_{1}+\dots+q_{m}\leq p_{1}+\dots+p_{n}$.
@@
 Let $t_{1},\dots,t_{n}$ be permission augmentations to $c_{1},\dots,c_{n}$, respectively.
-We say that $(c_{1},t_{1})\to_{\mathrm{df}}\dots\to_{\mathrm{df}}(c_{n},t_{n})$ is a consistent (permission-augmented) trace iff:
-  • $(c_{1},t_{1})$ is consistent
-  • $(c_{i},t_{i})\to_{\mathrm{df}}(c_{i+1},t_{i+1})$ is consistent for $i\in\{1,\dots,n\}$.
-The following lemma allows us to swap the order of execution of operators in two consistent steps.
-Lemma 8 ().
+We say that $(c_{1},t_{1})\to_{\mathrm{df}}\dots\to_{\mathrm{df}}(c_{n},t_{n})$ is a consistent (permission-augmented) trace iff for all $i\in\{1,\dots,n\}$, $(c_{i},t_{i})\to_{\mathrm{df}}(c_{i+1},t_{i+1})$ is consistent.
+We now show some facts about a consistently permission-augmented transition or trace, leading to the final Theorem 11. First, we show that two consistent steps commute.
+Lemma 9.
 Let $(c_{1},t_{1})\to^{o_{1}}_{\mathrm{df}}(c_{2},t_{2})\to^{o_{2}}_{\mathrm{df}}(c_{3},t_{3})$ be a consistent trace.
@@
 Then $c^{\prime}_{2}\to^{o_{1}}_{\mathrm{df}}c_{3}$ and there is a permission augmentation $t^{\prime}_{2}$ such that $(c_{1},t_{1})\to^{o_{2}}_{\mathrm{df}}(c^{\prime}_{2},t^{\prime}_{2})\to^{o_{1}}_{\mathrm{df}}(c_{3},t_{3})$ is consistent.
-Proof Sketech.
-If one of $o_{1}$ and $o_{2}$ is neither load nor store, or if both of them are loads, the result will always be consistent.
-If $o_{1}$ and $o_{2}$ are both stores, and they are the same node, then it must be the case that $c^{\prime}_{2}=c_{2}$ and we can use $t^{\prime}_{2}=t_{2}$.
-Otherwise, assume both $o_{1}$ and $o_{2}$ are load/store operators, and they are different nodes in the dataflow graph.
-Since they can both fire at $c_{1}$, and they are different nodes, their input values must be attached with disjoint permissions.
-Therefore, $o_{1}$ and $o_{2}$ have no dependencies, and they will give the same result no matter in which order we execute them.
+Proof Sketch.
+If one of $o_{1}$ and $o_{2}$ is neither load nor store, or if both of them are loads, they can trivially commute with the same result.
+Otherwise, one of $o_{1}$ and $o_{2}$ is a store operator (without loss of generality, assume $o_{1}$ is a store).
+For $i\in\{1,2\}$, let $l_{i}$ be the memory location accessed by $o_{i}$ and $p_{i}$ be the sum of $o_{i}$’s input permissions.
+Since they can both fire at $c_{1}$, they should have suitable permissions: $\texttt{write}\ l_{1}\leq p_{1}$ and $\texttt{read}\ l_{2}\leq p_{2}$.
+Since $p_{1}$ and $p_{2}$ are disjoint, we have $l_{1}\neq l_{2}$ Therefore, they are accessing different memory locations, and should give the same result no matter in which order we execute them.
 ∎
 Inductively applying this lemma, we can swap the execution of an operator all the way to the beginning if it can fire and has not fired yet.
-Lemma 9 ().
+Lemma 10.
 Let $(c_{1},t_{1})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}(c_{n+1},t_{n+1})$ be a consistent trace.
@@
 If we have the multiset inclusion $\{o^{\prime}_{1},\dots,o^{\prime}_{m}\}\subseteq\{o_{1},\dots,o_{n}\}$, then $c_{m+1}\to^{*}_{\mathrm{df}}c_{n+1}$.
-Practically, with this bounded confluence theorem, we can do bounded model checking for the confluence property: we perform symbolic execution to obtain a trace of symbolic configurations $c_{1}(\mathbf{x})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}(\mathbf{x})$, and then we can solve for a consistent augmentation $(t_{1},\dots,t_{n+1})$ for this trace using an SMT (z3, ) solver.
+Although we have omitted much of the technical details of the proof, we have formalized a machine-checkable proof of Theorem 11 in Verus (Lattuada et al., 2023), a verification language, and the proof can be found in our GitHub repository (Lin et al., 2024).
+Practically, with this bounded confluence theorem, we can do bounded model checking for the confluence property: we perform symbolic execution to obtain a trace of symbolic configurations $c_{1}(\mathbf{x})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}(\mathbf{x})$, and then we can solve for a consistent augmentation $(t_{1},\dots,t_{n+1})$ for this trace using an SMT solver.
 If there is a solution, then it follows from Theorem 11 that any other trace (firing operators only in the multiset $\{o_{1},\dots,o_{n}\}$, counting multiplicity) would still converge to $c_{n+1}(\mathbf{x})$ in the end.
+Algorithm 1 Full confluence checking algorithm (Section 5.2).
+1: Inputs:
+  • $c^{1}_{\mathrm{df}}(\mathbf{x}^{1}),\dots,c^{n}_{\mathrm{df}}(\mathbf{x}^{n})$: a list of dataflow cut points;
+  • $P(L,k)$: a permission algebra.
+2:If ConfluenceCheck succeeds, then the dataflow program is confluent.
+3:procedure ConfluenceCheck
+4:     $\psi\leftarrow\top$triangleright Permission constraints to be accumulated.
+5:     for $i\in\{1,\dots,n\}$ do $\triangleright$ Initialize permission variables at cut points
+[…]
+7:         for each channel $e$ and $m\in\{1,\dots,|\mathrm{Channel}(c^{i}_{\mathrm{df}}(\mathbf{x}^{i}),e)|\}$ do
+[…]
+[…]
+10:     for $i\in\{1,\dots,n\}$ do
+[…]
+12:         $\textbf{enqueue}(Q,(c^{i}_{\mathrm{df}}(\mathbf{x}^{i}),t_{i}))$
+13:         while $Q$ is non-empty do
+14:              $(c_{\mathrm{df}}(\mathbf{x}^{i}),t)\leftarrow\textbf{dequeue}(Q)$
+15:              for $(c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i}),t^{\prime})\in\textbf{step\textunderscore canonical}(c_{\mathrm{df}}(\mathbf{x}^{i}),t)$ do $\triangleright$ For each symbolic branch
+16:
+17:                  $\triangleright$ Constraints in Definition 7, where $o$ is the fired operator.
+18:                  $p_{\mathrm{in}}\leftarrow$ sum of $o$’s input permission variables in $t$
+19:                  $p_{\mathrm{out}}\leftarrow$ sum of $o$’s output permission variables in $t^{\prime}$
+20:                  $\psi\leftarrow\psi\wedge(p_{\mathrm{out}}\leq p_{\mathrm{in}})$
+21:                  if $o$ reads memory location $l\in L$ then
+22:                       $\psi\leftarrow\psi\wedge(\texttt{read}\ l\leq p_{\mathrm{in}})$
+23:                  else if $o$ writes to memory location $l\in L$ then
+24:                       $\psi\leftarrow\psi\wedge(\texttt{write}\ l\leq p_{\mathrm{in}})$
+25:
+26:                  for $j\in\{1,\dots,n\}$ do
+27:                       if $\exists\sigma:\mathbf{x}^{j}\to V(\mathbf{x}^{i}).\ c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i})=c^{j}_{\mathrm{df}}(\sigma)$ then $\triangleright$ Matched to a cut point
+28:                           for each channel $e$ and $m\in\{1,\dots,|\mathrm{Channel}(c^{j}_{\mathrm{df}}(\mathbf{x}^{j}),e)|\}$ do
+29:                                $\psi\leftarrow\psi\wedge(t^{\prime}(e,m)=t_{j}(e,m))$
+30:                           continue Line 15
+31:                  $\textbf{enqueue}(Q,(c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i}),t^{\prime}))$
+32:     succeed iff $\psi$ is satisfiable in $P(L,k)$
 5.2. Full Confluence Checking using Cut Points
 To extend the bounded confluence results in Section 5.1 to checking confluence for the entire dataflow program with unbounded traces, we need to finitely describe and check a consistent pattern of permission augmentations for all traces.
-For this purpose, we use the cut points selected in Section 4.
-Let $P^{1}_{\mathrm{df}}(\mathbf{x^{1}}),\dots,P^{n}_{\mathrm{df}}(\mathbf{x^{n}})$ be a list of cut points.
-During the bisimulation check in Section 4.2, we have verified that starting from any cut point $P^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, any branch of symbolic execution reaches an instance of another cut point $P^{j}_{\mathrm{df}}(\mathbf{x^{j}})$.
-A consequence of this check is that all the symbolic branches from all cut points “cover” any complete trace (traces from an initial configuration to a final configuration).
-In other words, any complete concrete trace from the initial configuration is the concatenation of instances of symbolic branches.
-Thus, if we are able to find a consistent permission augmentation for the trace of each symbolic branch during the bisimulation check, then any concrete trace in $\to_{\mathrm{df\mid canon}}$ would have a consistent permission augmentation, which implies that any other schedule will converge to $\to_{\mathrm{df\mid canon}}$, and the confluence of the entire dataflow program follows.
-Theorem 11 ().
-If all symbolic dataflow branches produced in Section 4.2 have consistent permission augmentations, then $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is confluent.
+For this purpose, we use the cut points selected in Section 4 and perform symbolic execution from them not only for symbolic values, but also for permissions.
+Let $P^{1},\dots,P^{n}$ be a list of cut points from the cut-simulation check (Section 4.3), where each $P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x}^{i}),c^{i}_{\mathrm{seq}}(\mathbf{y}^{i}))\mid\varphi^{i}(\mathbf{x}^{i},\mathbf{y}^{i})$ is a pair of symbolic dataflow and Seq configurations with an additional correspondence constraint.
+During the cut-simulation check in Section 4.3, we verify that starting from any dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x}^{i})$, any symbolic execution branch reaches an instance of another cut point $c^{j}_{\mathrm{df}}(\mathbf{x}^{j})$.
+A consequence is that the symbolic branches from all dataflow cut points “cover” any complete trace in the canonical schedule (Definition 3).
+In other words, any complete concrete trace in the canonical schedule is the concatenation of instances of symbolic branches.
+Thus, if we are able to find a consistent permission augmentation for the trace of each dataflow symbolic branch during the cut-simulation check, then any concrete trace in the canonical schedule would have a consistent permission augmentation, which implies that any other schedule will converge to the canonical schedule, and the confluence of the entire dataflow program follows.
+This full confluence check algorithm is procedurally described in Algorithm 1.
+The algorithm is parameterized with a list of dataflow cut points $c^{1}_{\mathrm{df}}(\mathbf{x}^{1}),\dots,c^{n}_{\mathrm{df}}(\mathbf{x}^{n})$ and a predetermined permission algebra $P(L,k)$.
+The algorithm first generates a fresh permission variable for each value present in the channels of the dataflow cut points.
+Then it performs symbolic execution from each cut point, carrying constraints for both symbolic values and their corresponding permission variables, until the symbolic branch is subsumed by another cut point.
+In the end, it checks for the satisfiability of the permission constraints.
+If they are satisfiable, the confluence of the full program follows.
+In more detail, the first loop at Line 5 assigns fresh permission variables to each value present in each cut point, representing an unknown concrete permission token to be solved for.
+We also add disjointness constraints at Line 9 to enforce that these permission augmentations are consistent (Definition 6) at the corresponding cut point configurations.
+Then in the main loop at Line 10, we symbolically execute from each dataflow cut point, trying to match each symbolic branch against other cut points.
+The step function step_canonical uses a deterministic canonical schedule (Definition 3) we constructed from Seq execution traces.
+For example, in the case of LLVM and RipTide, we use the canonical schedule described in Section 4.
+step_canonical is also instrumented to take a permission augmentation $t$ for the current configuration $c_{\mathrm{df}}(\mathbf{x}^{i})$, and output a modified permission augmentation $t^{\prime}$ for the stepped configuration $c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i})$, in which fresh variables are generated for all new output values in $c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i})$, and permissions corresponding to input values are removed.
+In each step, for each symbolic branch, we also accumulate permission constraints (Lines 17-24) to enforce that the augmented transition $(c_{\mathrm{df}}(\mathbf{x}^{i}),t)$ to $(c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i}),t^{\prime})$ is a consistent transition (Definition 7).
+At each symbolic step, we check if a symbolic branch is matched to another cut point $c^{j}_{\mathrm{df}}(\mathbf{x}^{j})$ (Line 27).
+This check is performed by finding a substitution $\sigma:\mathbf{y}^{j}\to\mathsf{Terms}(\mathbf{x}^{i})$ such that the current symbolic branch $c^{\prime}_{\mathrm{df}}(\mathbf{x}^{i})=c^{j}_{\mathrm{df}}(\sigma)$.
+If the symbolic branch matches the cut point, we add equality constraints (Line 29) to enforce that the permissions in $t^{\prime}$ should converge back to the original permissions $t_{j}$ at the matched $j$-th cut point.
+Note that the substitution $\sigma$ is not later used because we do not need to check correspondence constraints (Section 4.3).
+The main loop at Line 10 should terminate, because if the cut-simulation check (Section 4) passes, any branch from any dataflow cut point should eventually match another cut point.
+Finally, we check if the accumulated permission constraint $\psi$ is satisfiable in the permission algebra $P(L,k)$ via an SMT query.
+If $\psi$ is satisfiable, then any concrete trace (of unbounded length) in the canonical schedule has a consistent augmentation.
+Thus, by Theorem 11, any other possible schedule would eventually converge to the canonical schedule, and the dataflow program is confluent.
+If $\psi$ is not satisfiable, then there are different possibilities: 1) the dataflow program is not confluent and may have data races (which require manual inspection to confirm); 2) the choice of $k$ in the permission algebra $P(L,k)$ is not large enough; or 3) the confluence property is not provable using this method.
+5.3. Liveness in the Bounded-Channel Model
+In our current semantics for dataflow programs, channels between operators are modeled as unbounded queues.
+In real CGRA architectures, however, the channels have a fixed buffer size, and an operator may block if one of the output channels is full.
+As a result, a dataflow program in the bounded-channel model may have a deadlock that is not possible in the unbounded-channel model.
+For example, consider Figure 7, where operator A has two dataflow paths to B: one path is full, while the other is empty due to a Steer operator T discarding its inputs.
+This situation puts us in deadlock.
+Operator B cannot fire without tokens along both paths, while operator A cannot fire before more space is made along the full path.
+To verify that this scenario does not occur, we lift our cut-simulation and confluence results in the unbounded-channel model to the bounded-channel model via two observations:
+  • The symbolic branches in the dataflow cut point execution (Section 4.3) have finitely many intermediate symbolic configurations, so the maximum channel size is bounded by some integer $K$ in the canonical schedule.
+  • The confluence results in Section 5.1 and Section 5.2 are still valid in the bounded-channel model, which we have formalized and verified in Verus as well.
+Therefore, if the cut-simulation and the confluence checks pass in the unbounded-channel model, the same results hold in the bounded-channel model if channels have a buffer size of at least $K$.
+As a result, in the bounded-channel model, the dataflow program is still equivalent to the input imperative program, which guarantees the liveness property of the dataflow program that it will always make progress if the sequential imperative program is able to make progress.
+Furthermore, since the sequential program is deadlock-free, the dataflow program is also deadlock-free.
+Name LOC #OP #PC Sim. Conf. Description
+nn_vadd 5 15 244 0.02 0.17 Vector addition (RipTide)
+nn_norm 10 19 316 0.02 0.19 Neural network normalization (RipTide)
+nn_relu 11 19 844 0.03 0.45 Neural network ReLU layer (RipTide)
+smv 12 25 1103 0.05 2.27 Sparse-dense matrix-vector mult. (Pipestitch)
+dmv 10 31 1234 0.06 1.64 Dense-dense matrix-vector mult. (RipTide)
+Dither 19 31 2553 0.13 2.29 Dithering (Pipestitch)
+SpSlice 30 31 2616 $\times$ 0.13 7.94 Sparse matrix slicing (Pipestitch)
+nn_fc 23 36 2632 0.09 3.35 Neural network fully-connected layer (RipTide)
+SpMSpVd 45 40 3182 0.15 8.38 Sparse-sparse matrix-vector mult. (Pipestitch)
+bfs 48 48 3634 0.15 7.68 Breadth-first search (RipTide)
+dfs 45 49 3762 0.17 7.76 Depth-first search (RipTide)
+smm 22 49 3967 0.16 11.2 Sparse-dense matrix mult. (RipTide)
+nn_pool 32 52 8165 0.23 7.14 Neural network pooling layer (RipTide)
+dmm 16 56 4220 0.19 6.16 Dense-dense matrix mult. (RipTide)
+sconv 22 56 4272 0.2 10.37 Sparse convolution (RipTide)
+sort 23 57 6554 $\times$ 0.14 3.31 Radix sort (RipTide)
+SpMSpMd 39 60 7693 0.33 26.68 Sparse-sparse matrix mult. (Pipestitch)
+nn_conv 42 61 12154 0.2 15.32 Neural network convolution layer (RipTide)
+dconv 24 65 8134 0.25 11.51 Dense convolution (RipTide)
+fft 29 70 2690 0.12 5.77 Fast Fourier transform (RipTide)
+sha256 81 181 3548 $\times$ 0.28 12.6 SHA-256 hash
 Figure 5.
-Evaluation results on a benchmark of 20 programs.
-From left to right, the columns are names of the test cases (Name), lines of code in the original C program (LOC), numbers of dataflow operators (#OP), numbers of permission constraints (#PC), time spent for bisimulation check in seconds (Bisim.), time spent for confluence check in seconds (Conf.).
-All test cases are written originally in C and then compiled to LLVM via Clang (clang, ).
-The first 10 programs are benchmark programs from the RipTide work.
-The nn_* tests implement neural network inference.
-The last four are selected from a follow-up work to RipTide called Pipestitch (pipestitch, ).
-In the case of sort and SpSlice, the bisimulation check fails, because they contain instances of the same compiler bug.
-In all other cases, the bisimulation and confluence checks succeed.
-Name LOC #OP #PC Bisim. (sec) Conf. (sec)
-bfs 48 48 3634 0.45 7.79
-dconv 24 65 8134 0.96 11.87
-dfs 45 49 3762 0.46 8.51
-dmm 16 56 4220 0.71 6.31
-dmv 10 31 1234 0.22 1.69
-fft 29 70 2690 0.45 6.04
-sconv 22 56 4272 0.71 10.89
-smm 22 49 3967 0.53 12.08
-smv 12 25 1103 0.19 2.36
-sort 23 57 6554 $\times$ 0.64 3.62
-nn_conv 43 61 12154 0.78 15.65
-nn_fc 24 36 2632 0.31 3.43
-nn_norm 15 19 316 0.07 0.17
-nn_pool 33 52 8165 0.74 7.11
-nn_relu 16 19 844 0.11 0.44
-nn_vadd 9 15 244 0.06 0.15
-Dither 21 31 2553 0.32 2.36
-SpMSpMd 42 60 7693 0.93 28.45
-SpMSpVd 48 40 3182 0.4 10.22
-SpSlice 33 31 2616 $\times$ 0.34 7.4
+Evaluation results on 21 test cases sorted by the number of dataflow operators.
+From left to right, the columns are the name of the test case (Name), lines of code in the original C program (LOC), number of dataflow operators (#OP), number of permission constraints (#PC), time spent for simulation check in seconds (Sim.), and time spent for confluence check in seconds (Conf.).
+All test cases are originally written in C and then compiled to LLVM via Clang (LLVM, 2024a).
+Compiler bugs cause the simulation checks to fail in SpSlice, sort, and sha256.
+In all other cases, the simulation and confluence checks succeed.
+[…]
 Evaluation results.
-Figure 5.
-Evaluation results on a benchmark of 20 programs.
-From left to right, the columns are names of the test cases (Name), lines of code in the original C program (LOC), numbers of dataflow operators (#OP), numbers of permission constraints (#PC), time spent for bisimulation check in seconds (Bisim.), time spent for confluence check in seconds (Conf.).
-All test cases are written originally in C and then compiled to LLVM via Clang (clang, ).
-The first 10 programs are benchmark programs from the RipTide work.
-The nn_* tests implement neural network inference.
-The last four are selected from a follow-up work to RipTide called Pipestitch (pipestitch, ).
-In the case of sort and SpSlice, the bisimulation check fails, because they contain instances of the same compiler bug.
-In all other cases, the bisimulation and confluence checks succeed.
 6. Implementation and Evaluation
-We have implemented our translation validation technique in the tool WaveCert  [^2] Available on GitHub at https://github.com/secure-foundations/riptide-verification..
-We first implemented symbolic executors for (a subset of) LLVM and dataflow programs in Python (z3, ).
-Based on these, we implemented the bisimulation check in Section 4 and the confluence check in Section 5.
-WaveCert uses the Z3 SMT solver (z3, ) to discharge the verification conditions it generates.
-In the bisimulation check, WaveCert encodes the validity of the correspondence conditions at each cut point into Z3 queries.
-In the confluence check, WaveCert produces a list of permission constraints and checks if these constraints are satisfiable with respect to the finite permission algebra (Definition 2) by encoding each permission variable as a set of Boolean variables and the list of permission constraints as a SAT query.
-In order to evaluate WaveCert’s capability to certify compilation to real dataflow programs, we have applied WaveCert to a benchmark of 20 programs consisting of the benchmark programs in RipTide (riptide, ) and Pipestitch (pipestitch, ), as well as some programs implementing neural network inference.
+We have implemented our translation validation technique in the tool WaveCert targeting the RipTide compiler from LLVM programs to dataflow programs.
+It is publicly available in a GitHub repository (Lin et al., 2024).
+We implemented symbolic executors for (a subset of) LLVM and dataflow programs in Python.
+Based on these, we implemented the cut-simulation check described in Section 4 and the confluence check in Section 5.
+We also instrumented the RipTide compiler to output hints for constructing the canonical schedule (Section 4).
+WaveCert uses the Z3 SMT solver (De Moura and Bjørner, 2008) to discharge the verification conditions it generates.
+In the simulation check, WaveCert encodes feasibility of path conditions and validity of the correspondence conditions at each cut point into Z3 queries.
+In the confluence check, WaveCert produces a list of permission constraints and checks if these constraints are satisfiable with respect to the finite permission algebra (Definition 3) by encoding each permission variable as a set of Boolean variables and the list of permission constraints as an SMT query.
+In order to evaluate WaveCert’s capability to certify compilation to real dataflow programs, we have applied WaveCert to a benchmark of 21 programs consisting of the benchmark programs in RipTide (Gobieski et al., 2023) and Pipestitch (Serafin et al., 2023), as well as some programs implementing neural network inference and an implementation of SHA-256 (Conte, 2024).
 Figure 5 shows some statistics about the benchmark programs and how WaveCert performs.
-All tests are performed on a laptop with an M1 Max processor and 64 GiB of RAM.
-In all of the benchmark programs, the bisimulation check either succeeds or finds an actual bug (such as in sort and SpSlice, which we discuss further in Section 6.1).
-This shows that our heuristics for inferring dataflow cut points works well on the canonical schedule.
-Our bisimulation check generates verification conditions in quantifier-free theories of arrays and integers without non-linear arithmetic, so all bisimulation checks finish quickly within a second.
-The confluence check succeeds in all examples. For sort, we manually fixed the incorrect dataflow program produced by the compiler, after which the confluence check passes.
+All tests are performed on a laptop with an Apple M1 processor and 64 GiB of RAM.
+The time spent in the RipTide compiler (including hint generation) is less than 0.1 seconds for all of the benchmark programs, so it is omitted from the table.
+We now evaluate WaveCert in three aspects: generality of cut-point heuristics, performance, and size of the trusted computing base.
+Generality of the heuristics
+Both simulation and confluence checks pass for most of the benchmark programs.
+This shows that our heuristics for inferring dataflow cut points work well on the canonical schedule, and our confluence check is general enough for common compilation patterns.
+In test cases sort, SpSlice, and sha256, the simulation check fails and correctly reveals two compilation bugs (see Section 6.1 for more detail).
+After fixing these bugs, the simulation and confluence checks succeed for these two programs.
+Performance
+WaveCert spends most of its time on the confluence check, mainly consisting of checking satisfiability for permission constraints.
+The simulation check takes a relatively short amount of time, because the input LLVM program and the output dataflow program match well on the canonical schedule and thus the verification conditions are simple to solve.
+Overall, the verification time is within 15 seconds for most benchmark programs, and we believe it is feasible to enable WaveCert in real production scenarios.
+Trusted Computing Base
+The original RipTide compiler has ~21,000 lines of C++ code, excluding comments.
+The instrumented hint generation in the RipTide compiler is implemented in less than 50 lines of C++.
+Compared to the RipTide compiler, the trusted computing base of WaveCert has 3,121 lines of Python with three main components: ~800 lines of dataflow semantics; ~1,000 lines of LLVM semantics; and ~750 lines for simulation and confluence checks.
 Figure 6. Examples of two compiler bugs.
+[…]
 Examples of two compiler bugs.
-Figure 6. Examples of two compiler bugs.
+Figure 7. Deadlock example.
+[…]
+Deadlock example.
 6.1. Compiler Bugs
-During the testing of WaveCert, we found 7 bugs in the RipTide compiler.
-Three of these bugs were discovered on the test cases sort, SpSplice, and their variations; and the rest four bugs were found on smaller test cases that we manually constructed.
-A developer of the RipTide compiler has confirmed these issues.
-In the following, we discuss two examples of these bugs to show that the bisimulation check and the confluence check are both effective in catching compilation bugs.
+While testing WaveCert, we found 8 bugs in the RipTide compiler, confirmed by RipTide authors.
+Four of these bugs were discovered on the test cases sort, SpSplice, sha256, and their variations; and the remaining four bugs were found on test cases not included in the evaluation.
+We now discuss two representative bugs:
 Incorrect signed extension
@@
 A simplified LLVM program used to reproduce this issue is shown in the top left of Figure 7, and the output dataflow program is shown in the top right of Figure 7.
-The issue itself is straightforward: in the LLVM code, a 1-bit xor is done with a 1-bit constant 1, and then the result is zero-extended to a 32-bit integer.
-On the dataflow side, the dataflow compiler tries to convert all operations into 32-bits, the word width supported on the RipTide architecture.
+The LLVM code performs a 1-bit xor with a 1-bit constant 1, and it zero-extends the result to a 32-bit integer.
+On the dataflow side, the dataflow compiler converted all operations into 32-bits (word length of RipTide).
 However, during the process, the compiler assumes that all constants should be sign-extended.
 As a result, it extends the 1-bit 1 to a 32-bit -1, which is incorrect in this case.
-Our bisimulation check catches this bug when verifying the correspondence condition at the final cut point, where the memory configurations do not match.
+This difference causes our simulation check to fail, as the final memory states fail to match.
 Incorrect memory ordering with potential data races
-Our confluence check catches a compilation issue when compiling from the C program shown in the bottom left of Figure 6 to the dataflow program in the bottom right of Figure 6.
-In the C program, we have a nested loop.
-In each iteration of the outer loop, we first set A[i] = i, and then in the inner loop that follows, we copy A[i] into A[i + 1], …, A[n - 1].
+Our confluence check catches a compilation bug when compiling from the C program shown in the bottom left of Figure 7 to the dataflow program in the bottom right of Figure 7.
+The C program has a nested loop.
+In each iteration of the outer loop, it first sets A[i] = i, and then in the inner loop that follows, it copies A[i] into A[i + 1], …, A[n - 1].
 In the output dataflow program, the compiler allows the load and store operators in the inner loop to pipeline using the invariant operator (marked I) to repeatedly send the signal to enable load.
 However, this causes a data race, because in the inner loop, a load in the second iteration accesses the same memory location as the store in the first iteration.
-Our confluence check fails as the list of permission constraints we collected are unsatisfiable.
+The confluence check fails as the generated permission constraints are unsatisfiable.
 WaveCert also outputs an unsat core of 57 constraints, a subset of the 420 constraints generated, that causes the unsatisfiability.
-In particular, the unsat core includes two ordering constraints: read  A is contained in the disjoint union of input tokens of the inner loop load and write  A is contained in the disjoint union of input tokens of the inner loop store.
-This indicates that the main conflict stems from the inner loop load and store.
+In particular, the unsat core includes two constraints: read  A is in the input tokens of the inner loop load, and write  A is in the input tokens of the inner loop store.
+This indicates that the main conflict stems from these two operators.
 7. Limitations and Future Work
-Our translation validation technique still has some limitations, and we plan to address them in future work.
+Our translation validation technique still has some limitations to address in future work.
 Optimized dataflow graphs
@@
   • Array dependency analysis, which removes unnecessary orderings between loads and stores across iterations that access different indices of an array.
-The first two optimizations are not supported by our bisimulation check because we require a one-to-one correspondence between LLVM instructions and dataflow operators for the canonical schedule.
-In future work, we can relax this requirement and allow multiple LLVM instructions to map to one dataflow operator, and we then only fire the operator if all of the corresponding LLVM instructions have fired.
+The first two optimizations are not supported by our simulation check because we require a one-to-one correspondence between LLVM instructions and dataflow operators for the canonical schedule.
+Since these optimizations involve reasoning about the equivalence between dataflow programs, we believe that a modular approach is to use WaveCert to verify the input LLVM against the unoptimized dataflow program, and use another verification pass in future work to soundly optimize the dataflow program.
 The third optimization may lead to a dataflow program failing our confluence check.
@@
 This also requires a form of “dependent” permission tokens that can depend on the values in the channels.
-Liveness properties
-Our semantics of the dataflow program uses unbounded FIFO queues to model dataflow channels.
-In this model, we do not need to consider liveness issues such as deadlocks, because an operator can always fire once the input values are ready.
-However, this is not true for the actual hardware, which has a limit on the queue size.
-For future work, we plan to explore model checking (model-checking, ) techniques to check for deadlocks.
+Scalability of the confluence check
+WaveCert currently spends most of its time on the confluence check (Section 6).
+We believe this is primarily due to a naive encoding of permission constraints in SMT.
+For better performance in future work, one could design a custom solver for permission constraints to better utilize the structure of a permission algebra, or use a preprocessing step to simplify permission constraints.
 Lack of architecture-specific details
@@
 8. Related Work
-As a compiler verification technique, the translation validation approach taken in our work is a less labor-intensive alternative to full compiler verification, such as CompCert (compcert, ).
-It was first proposed by Pnueli et al.
-(pnueli-tv, ), and there are numerous following works with novel translation validation or program equivalence checking techniques for various languages and compilers (necula-tv, ; kundu-tv, ; data-driven-eq, ; tate-eq-saturation, ).
-For LLVM-related works on translation validation, we have Alive (alive, ; alive2, ), which performs bounded translation validation for internal optimizations in LLVM.
-LLVM-MD (llvm-md, ; denotational-tv, ) is another translation validation tool for LLVM optimizations.
-LLVM-MD translates the LLVM code into an intermediate representation called synchronous value graphs (SVG) (denotational-tv, ), which is a similar representation to a dataflow program.
-The difference is that the semantics of SVG is synchronous, and they also trust the compilation from LLVM to SVG, whereas in our case, such translation is exactly what we are trying to validate.
-Our formulation of cut-bisimulation in our bisimulation check is an adaptation from the work on Keq (keq, ), which aims to be a more general program equivalence checker parametric in the operational semantics of input/output languages.
-Our use of linear permission tokens is a dynamic variant of the fractional permissions proposed by Boyland (fractional-perm, ) and used in some variations of concurrent separation logic (concurrent-separation-logic, ).
-In contrast to separation logic frameworks such as Iris (iris, ), which are embedded in interactive proof assistants, our permission tokens are synthesized and checked fully automatically.
-There are also verification tools for classical dataflow languages such as Lustre (lustre, ; lustre-verified-compiler, ; lustre-smt-verification, ) and Esterel (esterel-verification, ).
-These languages have a synchronous semantics and do not interact with globally accessible memory, whereas dataflow programs emitted by CGRA compilers are highly asynchronous and make pervasive use of the global heap.
-In the context of term rewriting systems (term-rewriting, ), the standard approach for confluence checking is the use of critical pairs (knuth-critical-pairs, ; duran-confluence, ), which proves ground confluence in the case when the rewriting system is terminating.
-This algorithm is implemented in various systems, such as the Church-Rosser checker in Maude (maude, ).
+As a compiler verification technique, the translation validation approach taken in our work is a less labor-intensive alternative to full compiler verification, such as CompCert (Leroy, 2009).
+Translation validation was first proposed by Pnueli et al.
+(1998), and there are numerous following works with novel translation validation or program equivalence checking techniques for various languages and compilers (Necula, 2000; Kundu et al., 2009; Sharma et al., 2013; Tate et al., 2009).
+For LLVM-related works on translation validation, we have Alive (Lopes et al., 2015, 2021), which performs bounded translation validation for internal optimizations in LLVM.
+LLVM-MD (Tristan et al., 2011; Govereau, 2012) is another translation validation tool for LLVM optimizations.
+LLVM-MD translates the LLVM code into an intermediate representation called synchronous value graphs (SVG) (Govereau, 2012), which is a similar representation to a dataflow program.
+The difference is that the semantics of SVG is synchronous (in the sense that values are not buffered in the channels), and they also trust the compilation from LLVM to SVG, whereas in our case, such translation is exactly what we are trying to validate.
+Our formulation of cut-simulation in our simulation check is an adaptation from the work on Keq (Kasampalis et al., 2021), which aims to be a more general program equivalence checker parametric in the operational semantics of input/output languages.
+Our use of linear permission tokens is a dynamic variant of the fractional permissions proposed by Boyland (Boyland, 2013) and used in various separation logics (Brookes, 2006) to reason about read-only sharing of references.
+Tools based on separation logics such as Iris (Jung et al., 2018) and Viper (Müller et al., 2016) usually require manual annotations for permission passing in the pre-/post-conditions, whereas our permission tokens are synthesized fully automatically.
+Furthermore, separation logic is not a suitable formalism to directly reason about dataflow programs due to the lack of structured control-flow in a dataflow program and the asynchrony of dataflow operators.
+Various models of dataflow programming (Johnston et al., 2004) have been studied, such as computation graphs (Karp and Miller, 1966) and Kahn process networks (Gilles, 1974).
+An important aspect of these works is determinacy; i.e., these models produce a deterministic result regardless of the schedule of execution.
+There are also verification tools for classical dataflow languages such as Lustre (Halbwachs et al., 1991; Bourke et al., 2017; Hagen and Tinelli, 2008) and Esterel (Bouali, 1998).
+However, dataflow programs used by CGRA compilers often make use of shared global memory, which requires stronger techniques for guaranteeing determinacy.
+Process calculi such as CSP (Roscoe, 1997) and $\pi$-calculus (Milner, 1999) present an alternative formalism for message-passing concurrency.
+To manage the inherent complexity and nondeterminism in process calculi, session types can enforce confluence and deadlock-freedom by construction (Sangiorgi and Walker, 2001).
+However, similar to separation logic, session types require manual typing annotations that would be difficult to synthesize automatically, and we likely still need some form of fractional permissions to handle shared memory.
+Our strategy for analyzing nondeterministic dataflow programs is reminiscent of verification strategies for distributed systems (Bakst et al., 2017; Kragl et al., 2020; von Gleissenthall et al., 2019) which analyze message-passing protocols (e.g., Paxos (Lamport, 2001)) via reductions to sequential programs.
+However, the dataflow domain requires significantly different techniques: we verify equivalence properties with arbitrary specification programs, and use linear permissions to enable safe memory accesses in the presence of asynchrony.
+In the context of term rewriting systems (Baader and Nipkow, 1998), the standard approach for confluence checking is the use of critical pairs (Knuth and Bendix, 1970; Durán et al., 2020), which proves ground confluence in the case when the rewriting system is terminating.
+This algorithm is implemented in various systems, such as the Church-Rosser checker in Maude (Clavel et al., 2002).
 However, this technique is difficult to apply in our case, since the semantics of a dataflow program is neither terminating nor, in general, ground confluent.
 9. Conclusion
-In this work, we develop a translation validation technique for asynchronous dataflow programs with an application to the RipTide CGRA architecture.
-Our technique first checks the bisimulation of the input imperative program and the output dataflow program on a canonical schedule, and then we generalize the results to all possible schedules via linear permission tokens.
-Our work is an initial step towards integrating high assurance into new architectures from the start, rather than retrofitting them decades later.
+In this work, we develop a technique for translation validation between imperative source programs and asynchronous dataflow target programs, with an implementation for the RipTide CGRA architecture.
+Our technique simplifies analysis on dataflow by first verifying that the target program simulates the source program along a canonical schedule, and then using linear memory permissions to show that this canonical schedule is general.
+Our verification procedure ensures both functional correctness and liveness for the target dataflow program.
+In future work, we aim to support optimizations on dataflow programs (e.g., special-purpose operators for pipelining loops), and extending our verification methodology to concrete instantiations of CGRAs in hardware.